### v3

# Identifiability and Maximum Likelihood Estimation for System Identification of Networks of Dynamical Systems

Abstract
In this paper we investigate identifiability and maximum likelihood estimation for direct system identification of networks of dynamical systems. We provide necessary and sufficient conditions for network identifiability in terms of Gröbner bases. We show that the maximum likelihood approach is both consistent and efficient, which is in contrast to existing prediction error approaches. Moreover, our approach has wider applicability, i.e., it is applicable whenever network identifiability holds. Finally, we show that we can formulate the maximum likelihood problem without the use of a predictor, which is the key to numerically being able to solve it efficiently.
Keywords
System Identification, Networks, Dynamical Systems, Maximum likelihood estimation, Consistency, Efficiency.
I Introduction
System identification for networks of dynamical systems is an active area of research that has a long history. See [10] for a survey of early work. Two different ways of formulating the problem are the so-called indirect methods and the direct methods, see e.g. [1]. The indirect methods, e.g. [23, 14], model the closed loop transfer functions, and then try to recover the open loop transfer functions from the identified closed loop one. This is, in general, not numerically viable. Oftentimes the open loop transfer functions computed in this way will be unstable due to closed loop zeros that are estimated to be non minimum-phase. The direct methods, e.g. [22], instead directly estimate the open loop transfer functions from data obtained from closed loop experiments.
For both types of methods, it is highly relevant to develop conditions for network identifiability of the open loop transfer functions from closed loop data; see, e.g. [21] and the references therein. We will see that the network identifiability question can be efficiently addressed by considering it as a solvability question of a polynomial equation, which can be investigated by computing a Gröbner basis. We believe that this is a novel idea. The criterion is independent of what method is used for system identification.
Moreover, we propose a Maximum Likelihood (ML) approach for direct identification of the open loop transfer functions. We show that we have consistent estimates of the transfer functions from the inputs to the outputs of the individual systems in the network under fairly general assumptions. When this is the case, it is a trivial variation of [19] to show that the estimates are also efficient, i.e. the Cramér-Rao lower bound is achieved. Because of this, the ML formulation of the identification problem is highly desirable. In [22, 9] a consistent Prediction Error Method (PEM) for direct system identification of networks of systems is presented. As we will see, it cannot be applied to all cases for which network identifiability holds. Moreover, it is in many cases not efficient.
The ML formulation is in general difficult to solve numerically. This is because the optimization problem is non-convex in a much more complicated way than what is the case for system identification of e.g., ARMAX models. We will show how to circumvent this challenge by formulating the ML problem without explicit use of a predictor, which is the usual way to formulate the problem. The key to our equivalent formulation is to recognize that the problem can be looked at as a problem with latent variables or missing data, for which ML formulations have been developed in [13, 25].
The contributions of this paper can be summarized as:
  1. 1.
    Conditions for network identifiability in terms of Gröbner bases.
  2. 2.
    Conditions for consistency of ML system identification for networks.
  3. 3.
    Formulation of an ML problem for system identification as an unconstrained optimization problem.
Our contribution to consistency is a generalization of the work in [20] to general systems on innovation form and to networks of systems. In relation to [16] our work gives an alternative proof for the ergodic case, and it also details the conditions for the specific predictor we use in the context of networks. In relation to the work in [22, 9] we can obtain consistent estimates under less strict conditions. Moreover, our estimates are not only consistent but also efficient. Most of the results for Item 3 were presented in [8].
I-A Notation
We consider all vectors to be column matrices. With $\hat{e}_{i}$ we denote the $i$th standard basis vector. With $\mathbf{1}$ we denote the vector of all ones. If we have defined scalars or vectors $x_{k}$ for $1\leq k\leq n$, then we have also tacitly defined the vector $x=(x_{1},\ldots,x_{n})$. We denote by $A\otimes B$ the Kronecker product of $A$ with $B$. We use $X\succ$ and $X\succeq$ to denote positive definiteness and positive semidefiniteness of the symmetric matrix $X$, respectively. The notation $A\succ B$ is equivalent to $A-B\succ 0$, and similarly for positive semidefiniteness. We let boldface letters denote the Z-transform of signals. These signals can be scalar-valued, vector-valued or matrix-valued. We use the same notation for random variables as for realizations of random variables. With $\operatorname{\mathbb{E}}\!\left[\,X\,\right]$ is meant the expected value of the random variable $X$.
II Model of Network
We consider single-input single-output ARMAX models
$y_{k}^{i}+\sum_{j=1}^{n^{i}}a_{j}^{i}y_{k-j}^{i}=\sum_{j=1}^{n^{i}}b_{j}^{i}u_{k-j}^{i}+e_{k}^{i}+\sum_{j=1}^{n^{i}}c_{j}^{i}e_{k-j}^{i}$ (1)
for $1\leq i\leq M$, where $e^{i}_{k}$ is a sequence of independent zero mean Gaussian random variables. We assume that these sequences are independent of one another for different $i$ and that they have the variances $\lambda^{i}$. [^1] Notice that we do not need to assume that the $e_{k}$ is Gaussian in order to prove consistency. It is enough that it is a stationary stochastic process with zero mean and bounded second moment. However, we do use this assumption to formulate the maximum likelihood problem. We write $a^{i}=(a_{1}^{i},\ldots,a_{n^{i}}^{i})$ and we define $b^{i}$ and $c^{i}$ similarly. We also let $a=(a^{1},\ldots,a^{M})$ and we define $b$ and $c$ similarly. We let $n=\sum_{i=1}^{M}n^{i}$ and we define the polynomials
$\mathbf{A}^{i}(z)$ $=z^{n^{i}}+a_{1}^{i}z^{n^{i}-1}+\cdots+a_{n^{i}}$
$\mathbf{B}^{i}(z)$ $=b_{1}^{i}z^{n^{i}-1}+\cdots+b_{n^{i}}$
$\mathbf{C}^{i}(z)$ $=z^{n^{i}}+c_{1}^{i}z^{n^{i}-1}+\cdots+c_{n^{i}}$
in the complex variable $z$. We make the standing assumption that
  1. A1.
    $\mathbf{A}^{i}(z)$ and $\mathbf{B}^{i}(z)$ are coprime.
We also define the transfer functions
$\mathbf{G}^{i}(z)=\frac{\mathbf{B}^{i}(z)}{\mathbf{A}^{i}(z)},\quad\mathbf{H}^{i}(z)=\frac{\mathbf{C}^{i}(z)}{\mathbf{A}^{i}(z)}.$
and $\mathbf{G}(z)=\operatorname{blkdiag}\left(\mathbf{G}^{1}(z),\ldots,\mathbf{G}^{M}(z)\right)$ and $\mathbf{H}(z)$ similarly. A model is defined as the vector $\mathbf{M}(z)=(\mathbf{G}(z),\mathbf{H}(z))$.
For each ARMAX model we introduce an observer form realization
$\xi_{k+1}^{i}$ $=F^{i}\xi_{k}^{i}+b^{i}u^{i}_{k}+(c^{i}-a^{i})e_{k}^{i}$
$y^{i}_{k}$ $=H^{i}\xi^{i}_{k}+e_{k}^{i},$
where $F^{i}=U-a^{i}\hat{e}_{1}^{T}$, $U$ is a shift matrix with ones on the first super-diagonal, and where $H^{i}=\hat{e}_{1}^{T}$. Let $\xi_{k}=(\xi^{1}_{k},\ldots,\xi^{M}_{k})$, and define $u_{k}$, $y_{k}$ and $e_{k}$ similarly. Then with
$F$ $=\operatorname{blkdiag}(F^{1},\ldots,F^{M})$
$B$ $=\operatorname{blkdiag}(b^{1},\dots,b^{M})$
$C$ $=\operatorname{blkdiag}(c^{1}-a^{1},\dots,c^{M}-a^{M})$
$H$ $=\operatorname{blkdiag}(H^{1},\dots,H^{M}),$
it holds that
$\xi_{k+1}$ $=F\xi_{k}+Bu_{k}+Ce_{k}$
$y_{k}$ $=H\xi_{k}+e_{k}.$
Let $\Upsilon\in\mathbb{R}^{M\times M}$ and $\Omega\in\mathbb{R}^{M\times m}$ be zero-one matrices. The closed-loop system [^2] We assume that the closed loop system is well-posed. for $u_{k}=\Upsilon y_{k}+\Omega r_{k}$ is given by
$\xi_{k+1}$ $=F_{c}\xi_{k}+G_{r}r_{k}+G_{e}e_{k}$
$\begin{bmatrix}y_{k}\\ u_{k}\end{bmatrix}$ $=H_{c}\xi_{k}+J_{r}r_{k}+J_{e}e_{k},$
where $F_{c}=F+B\Upsilon H,\;G_{r}=B\Omega,\;G_{e}=C+B\Upsilon$ and
$H_{c}=\begin{bmatrix}I\\ \Upsilon\end{bmatrix}H,\quad J_{r}=\begin{bmatrix}0\\ \Omega\end{bmatrix},\quad J_{e}=\begin{bmatrix}I\\ \Upsilon\end{bmatrix}$.
We then define the observed output as
$x_{o,k}=T_{o}\begin{bmatrix}y_{k}\\ u_{k}\end{bmatrix}$
and with $H_{o}=T_{o}H_{c}$, $J_{ro}=T_{o}J_{r}$ and $J_{eo}=T_{o}J_{e}$ it holds that
$\xi_{k+1}$ $=F_{c}\xi_{k}+G_{r}r_{k}+G_{e}e_{k}$ (2)
$x_{o,k}$ $=H_{o}\xi_{k}+J_{ro}r_{k}+J_{eo}e_{k}.$
We remark that $e_{k}$ has covariance $\Sigma_{e}=\operatorname{diag}(\lambda^{1},\ldots,\lambda^{M})$, which we tacitly assume is positive definite throughout this paper. We denote with $\lambda$ the vector of all $\lambda^{i}$. The matrix $T_{o}$ is full row rank, and its $p$ rows are standard basis vectors.
We denote the closed loop transfer functions from $\mathbf{r}$ and $\mathbf{e}$ to $\mathbf{x}_{o}$ with $\mathbf{G}_{c}(z)$ and $\mathbf{H}_{c}(z)$, respectively. We also define $\mathbf{M}_{c}(z)=(\mathbf{G}_{c}(z),\mathbf{H}_{c}(z))$. We define the parameter vector $\theta=(a,b,c,\lambda)$. We denote with $\Theta_{o}$ the set of parameters $\theta$ such that $\mathbf{M}_{c}(z)$ have all poles strictly inside the unit circle. We also define the compact set $\Theta\subset\Theta_{o}$ such that it contains the parameter $\theta_{0}$ which represents the closed loop system from which we will collect the data $x_{o}$. It is over this set we will later on optimize a likelihood function. The model set $\mathcal{M}$ is the set of all models $\mathbf{M}_{c}(z)$ such that $\theta\in\Theta$. Sometimes we will explicitly write e.g. $\mathbf{M}_{c}(z;\theta)$ to emphasize the parameter dependence $\theta$ of the transfer function. This also goes for all other variables. We will sometimes tacitly assume that $\theta\in\Theta$.
III Predictor
We equivalently describe the observations using the innovation form. To this end, we introduce the equation
$\begin{bmatrix}I&K\\ 0&I\end{bmatrix}\begin{bmatrix}\Sigma&0\\ 0&\Sigma_{\epsilon}\end{bmatrix}\begin{bmatrix}I&K\\ 0&I\end{bmatrix}^{T}$ (3)
$\qquad=\begin{bmatrix}F_{c}&G_{e}\\ H_{o}&J_{eo}\end{bmatrix}\begin{bmatrix}\Sigma&0\\ 0&\Sigma_{e}\end{bmatrix}\begin{bmatrix}F_{c}&G_{e}\\ H_{o}&J_{eo}\end{bmatrix}^{T}$
in the variables $\Sigma$, $\Sigma_{\epsilon}$ and $K$. Solving for $K$ in the (1,2)-block and substituting into the (1,1)-block gives the Riccati equation for the variable $\Sigma$. The (3,3)-block defines $\Sigma_{\epsilon}=H_{o}\Sigma H_{o}^{T}+J_{eo}\Sigma_{e}J_{eo}^{T}$ which is the covariance for the innovations. The innovation form is then given by
$\hat{\xi}_{k+1}$ $=F_{c}\hat{\xi}_{k}+G_{r}r_{k}+K\epsilon_{k}$ (4)
$x_{o,k}$ $=H_{o}\hat{\xi}_{k}+J_{ro}r_{k}+\epsilon_{k},$
where $\epsilon_{k}$ are the innovations, which are such that $\epsilon_{j}$ is independent of $\epsilon_{k}$ whenever $j\neq k$. Moreover, it is zero mean and Gaussian with covariance $\Sigma_{\epsilon}$, as already mentioned. The innovation form also defines the Kalman predictor
$\hat{x}_{o,k}=H_{o}\hat{\xi}_{k}+J_{ro}r_{k},$
and we have that the prediction error is the innovation $\epsilon_{k}=x_{o,k}-\hat{x}_{o,k}=x_{o,k}-H_{o}\hat{\xi}_{k}-J_{ro}r_{k}$. From this we realize that the predictor is governed by the equations
$\hat{\xi}_{k+1}$ $=F_{c}\hat{\xi}_{k}+G_{r}r_{k}+K\left(x_{o,k}-H_{o}\hat{\xi}_{k}-J_{ro}r_{k}\right)$ (5)
$\hat{x}_{o,k}$ $=H_{o}\hat{\xi}_{k}+J_{ro}r_{k},$
which is the Kalman filter. Therefore, the predictor only depends on $r_{k}$ and on what we observe, i.e., $x_{o,k}$. The transfer function from $\mathbf{x}_{o}$ and $\mathbf{r}$ to $\hat{\mathbf{x}}_{0}$ is given by $\mathbf{W}(z)=(\mathbf{W}_{o}(z),\mathbf{W}_{r}(z))$, where
$\mathbf{W}_{o}(z)$ $=H_{o}(zI-F_{c}+KH_{o})^{-1}K$ (6)
$\mathbf{W}_{r}(z)$ $=H_{o}(zI-F_{c}+KH_{o})^{-1}(G_{r}-KJ_{ro})+J_{ro}.$
Lemma 1
Assume that $\theta\in\Theta$. There is a unique solution $(\Sigma,K,\Sigma_{\epsilon})$ to (3) such that the matrix $F_{c}-KH_{o}$ has all its eigenvalues strictly inside the unit disc, and such that $\Sigma_{\epsilon}$ is positive definite if and only if
  • A2.
    $J_{eo}$ has full row rank,
  • A3.
    $\mathbf{C}^{i}(z)\neq 0$ for $|z|=1$ and for all $i$.
Proof:
There is a unique solution $(\Sigma,K,\Sigma_{\epsilon})$ to (3) such that the matrix $F_{c}-KH_{o}$ has all its eigenvalues strictly inside the unit disc, and such that $\Sigma_{\epsilon}$ is positive definite if and only if $(F_{c},H_{o})$ is detectable and the pencil
$\mathbf{P}(z)=\begin{bmatrix}-zI+F_{c}&G_{e}\\ H_{o}&J_{eo}\end{bmatrix}$
has full row rank for $|z|=1$, see [12, Theorem 2]. Detectability is trivially fulfilled, since $F_{c}$ has all eigenvalues strictly inside the unit circle by the assumption that $\theta\in\Theta$. We can write the pencil as
$\mathbf{P}(z)=\begin{bmatrix}I&0\\ 0&J_{eo}\end{bmatrix}\begin{bmatrix}I&B\Upsilon\\ 0&I\end{bmatrix}\begin{bmatrix}-zI+F&C\\ H&I\end{bmatrix}$.
The first matrix has full row rank if and only if $J_{eo}$ has full row rank. The middle matrix is clearly invertible. The right matrix can after permutations of rows and columns be seen to be a block diagonal matrix with blocks
$\mathbf{Q}^{i}(z)$ $=\begin{bmatrix}-zI+F^{i}&c^{i}-a^{i}\\ H^{i}&1\end{bmatrix}$
$=\begin{bmatrix}-z-a_{1}^{i}&1&0&c_{1}^{i}-a_{1}^{i}\\ -a_{2}^{i}&-z&1&c_{2}^{i}-a_{2}^{i}\\ -a_{3}^{i}&0&-z&c_{3}^{i}-a_{3}^{i}\\ 1&0&0&1\end{bmatrix}$
for the case of $n^{i}=3$. We have with
$\mathbf{U}(z)=\begin{bmatrix}z^{2}&z&1&0\\ 0&1&0&0\\ 1&0&0&0\\ 0&0&0&1\end{bmatrix},\quad\mathbf{V}(z)=\begin{bmatrix}1&0&0&-1\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}$
that
$\mathbf{U}(z)\mathbf{Q}^{i}(z)\mathbf{V}(z)=\begin{bmatrix}-\mathbf{A}^{i}(z)&0&0&\mathbf{C}^{i}(z)\\ -a_{2}^{i}&-z&1&c_{2}^{i}\\ -z-a_{1}^{i}&1&0&z+c_{1}^{i}\\ 1&0&0&0\end{bmatrix}$.
We can, after this, multiply with further unitary matrices from the left and the right to see that the matrix will be invertible for $|z|=1$ if and only if the polynomial $\mathbf{C}^{i}(z)$ is not equal to zero for $|z|=1$. ∎
Remark 1
Assumption A2 can always be fulfilled by removing redundant observed signals. To see this, let $U$ be an invertible matrix such that
$\bar{J}_{eo}=\begin{bmatrix}\bar{J}_{eo1}\\ 0\end{bmatrix}=UJ_{eo},\quad\bar{J}_{ro}=\begin{bmatrix}\bar{J}_{ro1}\\ \bar{J}_{ro2}\end{bmatrix}=UJ_{ro}$
with $\bar{J}_{eo1}$ full row rank, and where the rows of $\bar{J}_{eo1}$ are a subset of the rows of $J_{eo}$. Then
$\bar{x}_{o,k}=\begin{bmatrix}\bar{x}_{o1,k}\\ \bar{x}_{o2,k}\end{bmatrix}=Ux_{o,k}=\begin{bmatrix}\bar{J}_{eo1}(H\xi_{k}+e_{k})+\bar{J}_{ro1}r_{k}\\ \bar{J}_{ro2}r_{k}\end{bmatrix}$.
The observation $\bar{x}_{o2,k}$ is not going to affect the prediction, and hence can be discarded. This holds since after transformation with $U$ there will be rows in the parenthesis in (5) that are zero. Hence, a $K$ with fewer columns is sufficient.
We will need equivalent formulations of the algebraic Riccati equation. To this end, let
$\begin{bmatrix}Q&S\\ S^{T}&R\end{bmatrix}=\begin{bmatrix}G_{e}\\ J_{eo}\end{bmatrix}\Sigma_{e}\begin{bmatrix}G_{e}\\ J_{eo}\end{bmatrix}^{T}.$
We then have that (3) is equivalent to
$\begin{bmatrix}I&K\\ 0&I\end{bmatrix}\begin{bmatrix}\Sigma&0\\ 0&\Sigma_{\epsilon}\end{bmatrix}\begin{bmatrix}I&K\\ 0&I\end{bmatrix}^{T}=\begin{bmatrix}F_{c}\\ H_{o}\end{bmatrix}\Sigma\begin{bmatrix}F_{c}\\ H_{o}\end{bmatrix}^{T}+\begin{bmatrix}Q&S\\ S^{T}&R\end{bmatrix}$. (7)
Multiply this equation with $\begin{bmatrix}I&-R^{-1}S\\ 0&I\end{bmatrix}$ from the left and its transpose from the right to obtain
$\Sigma+\bar{K}\Sigma_{\epsilon}\bar{K}^{T}$ $=\bar{F}_{c}\Sigma\bar{F}_{c}^{T}+\bar{Q}$ (8)
$\bar{K}\Sigma_{\epsilon}$ $=\bar{F}_{c}\Sigma H_{o}^{T}$
$\Sigma_{\epsilon}$ $=H_{o}\Sigma H_{o}^{T}+R,$
where $\bar{F}_{c}=F_{c}-SR^{-1}H_{o}$, $\bar{Q}=Q-SR^{-1}S^{T}$ and $\bar{K}=K-SR^{-1}$. By solving for $\bar{K}$ in the second equation and substitution into the first equation we obtain
$\Sigma=\bar{F}_{c}\Sigma\bar{F}_{c}^{T}+\bar{Q}-\bar{F}_{c}\Sigma H_{o}^{T}(H_{o}\Sigma H_{o}^{T}+R)^{-1}H_{o}\Sigma\bar{F}_{c}^{T}.$ (9)
This is an algebraic Riccati equation for the case when there is no cross covariance. There are cases for which there is a trivial solution to the algebraic Riccati equation.
Theorem 2
Assume that $\theta\in\Theta$, that Assumption A2 holds and that $\mathbf{C}^{i}(z)\neq 0$ for $|z|\geq 1$ and for all $i$. If
$\Sigma_{e}-\Sigma_{e}J_{eo}^{T}R^{-1}J_{eo}\Sigma_{e}=0,$ (10)
then $(\Sigma,K,\Sigma_{\epsilon})=(0,SR^{-1},R)$ is a solution to (3) such that $F_{c}-KH_{o}$ has all eigenvalues strictly inside the unit circle.
Proof:
Under the assumption we have that $\bar{Q}=Q-SR^{-1}S^{T}=0$, from which is straightforward to see that $(\Sigma,K,\Sigma_{\epsilon})=(0,SR^{-1},R)$ satisfies the equivalent formulation of the algebraic Riccati equation in (8). It holds that
$\bar{F}_{c}$ $=F_{c}-KH_{o}=F+B\Upsilon H-SR^{-1}H_{o}$
$=F+\left(B\Upsilon-G_{e}\Sigma_{e}J_{eo}^{T}R^{-1}J_{eo}\right)H$
$=F+(B\Upsilon-B\Upsilon-C)H=F-CH,$
where we have used the definition of $G_{e}$ and the fact that $J_{eo}^{T}R^{-1}J_{eo}=\Sigma_{e}^{-1}$ by our assumption. We see that $\bar{F}_{c}$ is a block diagonal matrix with blocks $U-a^{i}\hat{e}_{1}^{T}-(c^{i}-a^{i})\hat{e}_{1}^{T}=U-c^{i}\hat{e}_{1}^{T}$, which are on observer form. From our assumption on $\mathbf{C}^{i}$ it now follows that all blocks have their eigenvalues strictly inside the unit circle. ∎
Remark 2
One example of when (10) is fulfilled is when $T_{o}=\begin{bmatrix}I&0\end{bmatrix}$, that is when $x_{o,k}=y_{k}$.
From the definitions of the matrices it is straightforward to prove that the following lemma holds.
Lemma 3
It holds that $\Upsilon\Sigma_{e}J_{eo}^{T}R^{-1}$ is a zero-one matrix for any $\Sigma_{e}\succ 0$ irrespective of what $\lambda$ is.
For OE models we have the following theorem.
Theorem 4
Assume that $\theta\in\Theta$, that Assumption A2 holds, that $\mathbf{A}^{i}(z)\neq 0$ for $|z|\geq 1$ for all $i$, and that $a=c$. If
$\Upsilon\left(I-J_{eo}^{T}\left(J_{eo}J_{eo}^{T}\right)^{-1}J_{eo}\right)=0$ (11)
then $(\Sigma,K,\Sigma_{\epsilon})=(0,SR^{-1},R)$ is a solution to (3) such that $F_{c}-KH_{o}$ has all eigenvalues inside the unit circle.
Proof:
We have that
$\bar{Q}$ $=B\Upsilon\left(\Sigma_{e}-\Sigma_{e}J_{eo}^{T}R^{-1}J_{eo}\Sigma_{e}\right)\Upsilon^{T}B^{T}$
$=B\left(\Upsilon-\Upsilon\Sigma_{e}J_{eo}^{T}R^{-1}J_{eo}\right)\Sigma_{e}\Upsilon^{T}B^{T}.$
We have by Lemma 3 that the matrix in the parenthesis is a zero-one matrix. Hence, it is sufficient to let $\Sigma_{e}=I$, and then it follows from our assumption that $\bar{Q}=0$. As in the proof of the previous theorem, we obtain that $(\Sigma,K,\Sigma_{\epsilon})=(0,SR^{-1},R)$ satisfies the algebraic Riccati equation. We have
$F_{c}-KH_{o}$ $=F+B\Upsilon H-B\Upsilon\Sigma_{e}J_{eo}^{T}R^{-1}J_{eo}H$
$=F+B\left(\Upsilon-\Upsilon\Sigma_{e}J_{eo}^{T}R^{-1}J_{eo}\right)H=F$
by the same reasoning as above. From the assumptions on $\mathbf{A}^{i}(z)$ it follows that $F$ has all eigenvalues strictly inside the unit circle. ∎
Corollary 4.1
Under the assumptions in the theorem, it follows that the predictor transfer functions in (6) are given by
$\mathbf{W}_{o}(z)$ $=H_{o}(zI-F)^{-1}K$ (12)
$\mathbf{W}_{r}(z)$ $=H_{o}(zI-F)^{-1}(G_{r}-KJ_{ro})+J_{ro}$
where $K$ and $G_{r}-KJ_{ro}$ are equal to $B$ times zero-one matrices.
Proof:
We have $F_{c}-KH_{o}=F$ from the proof of the theorem. Moreover, $K=B\Upsilon\Sigma_{e}J_{eo}^{T}R^{-1}$ and
$G_{r}-KJ_{ro}=B\left(\Omega-\Upsilon\Sigma_{e}J_{eo}^{T}R^{-1}J_{ro}\right).$
The result then follows from Lemma 3. ∎
Remark 3
We notice that the predictor transfer functions are $\theta$-independent affine zero-one transformations of the open loop transfer function $\mathbf{G}(z)$. In case (11) does not hold, then this is not true, which follows from the fact that then $\Sigma\neq 0$.
Lemma 5
Under assumptions A2–A3, it holds that the solution $(\Sigma,K,\Sigma_{\epsilon})$ to (3) is a continuously differentiable function of $\theta\in\Theta_{o}$.
Proof:
The results follow from the proof of Theorem 4 in [12]. ∎
IV Informative Signals
The signals $(x_{o},r)$ are said to be informative enough if for any two models $\mathbf{M}_{c}(z;\theta_{1})$ and $\mathbf{M}_{c}(z;\theta_{2})$ in $\mathcal{M}$, it holds that
$\operatorname{\mathbb{E}}\!\left[\,\left(\hat{x}_{o,k}(\theta_{1})-\hat{x}_{o,k}(\theta_{2})\right)\left(\hat{x}_{o,k}(\theta_{1})-\hat{x}_{o,k}(\theta_{2})\right)^{T}\,\right]=0$ (13)
implies that $\mathbf{W}\left(e^{i\omega};\theta_{1}\right)=\mathbf{W}\left(e^{i\omega};\theta_{2}\right)$ for almost all $\omega$. The expectation is here with respect to the innovation $\epsilon_{k}(\theta_{0})$ as well as the reference value $r_{k}$, which is deterministic but assumed to be what is called quasi-stationary. Hence the expectation should be interpreted as an infinite time-average. For a more thorough discussion see [17, Section 2.3]. We will use superscript $i$ in this section to distinguish between different $\theta_{i}$. With boldface letters denoting Z-transforms, it holds that
$\Delta\boldsymbol{\epsilon}=\boldsymbol{\epsilon}^{1}-\boldsymbol{\epsilon}^{2}=\hat{\mathbf{x}}_{o}^{2}-\hat{\mathbf{x}}_{o}^{2}=\Delta\mathbf{W}_{r}\mathbf{r}+\Delta\mathbf{W}_{o}\mathbf{x}_{o}$
where $\Delta\mathbf{W}_{r}=\mathbf{W}_{r}^{2}-\mathbf{W}_{r}^{1}$ and $\Delta\mathbf{W}_{o}=\mathbf{W}_{o}^{2}-\mathbf{W}_{o}^{1}$. The observed signal $x_{o}$ is obtained from the model, which is indexed with zero, and hence we have
$\hat{\mathbf{x}}_{o}^{0}=\mathbf{W}_{r}^{0}\mathbf{r}+\mathbf{W}_{o}^{0}\mathbf{x}_{o},\quad\boldsymbol{\epsilon}^{0}=\mathbf{x}_{o}-\hat{\mathbf{x}}_{o}^{0}.$
Solving for $\mathbf{x}_{o}$ in these two equations and substituting into the previous equation results in
$\Delta\boldsymbol{\epsilon}$ $=\left(\Delta\mathbf{W}_{r}+\Delta\mathbf{W}_{o}(I-\mathbf{W}_{o}^{0})^{-1}\mathbf{W}_{r}^{0}\right)\mathbf{r}$
$\quad+\Delta\mathbf{W}_{o}\left((I-\mathbf{W}_{o}^{0})^{-1}\mathbf{W}_{o}^{0}+I\right)\boldsymbol{\epsilon}^{0}.$
First we notice that $(I-\mathbf{W}_{o}^{0})^{-1}\mathbf{W}_{o}^{0}+I=(I-\mathbf{W}_{o}^{0})^{-1}$. Then we notice that a realization of $(I-\mathbf{W}_{o}^{0})^{-1}$ is given by $(F_{c},K,H_{o},I)$, where we know that $F_{c}$ has all eigenvalues strictly inside the unit circle. Hence all transfer functions above have their poles strictly inside the unit circle. Let us write
$\Delta\boldsymbol{\epsilon}=\bar{\mathbf{W}}_{r}r+\bar{\mathbf{W}}_{o}\boldsymbol{\epsilon}^{0}$
where
$\bar{\mathbf{W}}_{r}$ $=\left(\Delta\mathbf{W}_{r}+\Delta\mathbf{W}_{o}(I-\mathbf{W}_{o}^{0})^{-1}\mathbf{W}_{r}^{0}\right)$
$\bar{\mathbf{W}}_{o}$ $=\Delta\mathbf{W}_{o}(I-\mathbf{W}_{o}^{0})^{-1}.$
The left hand side of (13) can, because of stationarity, be expressed as
$\begin{bmatrix}\bar{\mathbf{W}}_{r}\left(e^{i\omega}\right)&\bar{\mathbf{W}}_{o}\left(e^{i\omega}\right)\end{bmatrix}\!\!\begin{bmatrix}\Phi_{r}(\omega)&\Phi_{r,\epsilon}(\omega)\\ \Phi_{r,\epsilon}(-\omega)^{T}&\Sigma_{\epsilon}\end{bmatrix}\!\!\begin{bmatrix}\bar{\mathbf{W}}_{r}\left(e^{-i\omega}\right)^{T}\\ \bar{\mathbf{W}}_{o}\left(e^{-i\omega}\right)^{T}\end{bmatrix}$
where $\Phi_{r}(\omega)$ is the spectrum of $r$, and $\Phi_{r,\epsilon}(\omega)$ is the cross spectrum of $r$ and $\epsilon$. Since $\Sigma_{\epsilon}$ is positive definite, it holds that (13) implies that $\bar{\mathbf{W}}_{o}\left(e^{i\omega}\right)=0$ for almost all $\omega$. We notice that
$\begin{bmatrix}I&-K\\ 0&I\end{bmatrix}\begin{bmatrix}-zI+F_{c}&K\\ H_{o}&I\end{bmatrix}=\begin{bmatrix}-zI+F_{c}-KH_{o}&0\\ H_{o}&I\end{bmatrix}$,
and since by Lemma 1 it holds that $F_{c}-KH_{o}$ has all eigenvalues strictly inside the unit disc, it holds that $(I-\mathbf{W}_{o}^{0})^{-1}$ has no transmission zeros on the unit circle. We can therefore conclude that (13) implies that $\Delta\mathbf{W}_{o}\left(e^{i\omega}\right)=0$ for almost all $\omega$. From this it follows that $\bar{\mathbf{W}}_{r}=\Delta\mathbf{W}_{r}$, and hence we see that if $\Phi_{r}(\omega)$ is positive definite, then (13) implies that $\Delta\mathbf{W}_{r}\left(e^{i\omega}\right)=0$ for almost all $\omega$. Because of this, we introduce the assumption
  1. A4.
    $\Phi_{r}(\omega)\succ 0,\quad\forall\omega$,
which together with assumptions A2 and A3 will guarantee that the signals $(x_{o},r)$ are informative enough. We summarize this result in a lemma.
Lemma 6
Assume that A2–A4 holds. Then the signals $(x_{o},r)$ are informative enough, i.e. for any two models $\mathbf{M}_{c}(z;\theta_{1})$ and $\mathbf{M}_{c}(z;\theta_{2})$ in $\mathcal{M}$, it holds that (13) implies that $\mathbf{W}\left(e^{i\omega};\theta_{1}\right)=\mathbf{W}\left(e^{i\omega};\theta_{2}\right)$ for almost all $\omega$.
V Identifiability
We will now show that it is possible to obtain the innovation form from $\mathbf{W}$. Let
$\mathbf{G}_{o}(z)=H_{o}(zI-F_{c})^{-1}K.$ (14)
Applying the Z-transform to the innovation form results in
$\hat{\mathbf{x}}_{o}=\mathbf{G}_{c}\mathbf{r}+\mathbf{G}_{o}\boldsymbol{\epsilon},\quad\mathbf{x}_{o}=\hat{\mathbf{x}}_{o}+\boldsymbol{\epsilon}.$
We solve for $\hat{\mathbf{x}}_{o}$ in these equations and obtain that
$\hat{\mathbf{x}}_{o}=(I+\mathbf{G}_{o})^{-1}\mathbf{G}_{c}\mathbf{r}+(I+\mathbf{G}_{o})^{-1}\mathbf{G}_{o}\mathbf{x}_{o},$
and hence it must hold that
$\mathbf{W}_{o}=(I+\mathbf{G}_{o})^{-1}\mathbf{G}_{o},\quad\mathbf{W}_{r}=(I+\mathbf{G}_{o})^{-1}\mathbf{G}_{c}.$
We can now solve for $\mathbf{G}_{o}$ and $\mathbf{G}_{c}$ in the above equations, which results in
$\mathbf{G}_{o}$ $=\mathbf{W}_{o}(I-\mathbf{W}_{o})^{-1}$ (15)
$\mathbf{G}_{c}$ $=\left(I+\mathbf{W}_{o}(I-\mathbf{W}_{o})^{-1}\right)\mathbf{W}_{r}.$
From this, we realize that we can recover the innovation form for the closed loop system from $\mathbf{W}$. We summarize this result in a lemma.
Lemma 7
Given the predictor transfer functions $(\mathbf{W}_{o},\mathbf{W}_{r})$, it is possible to recover the innovation form transfer functions $(\mathbf{G}_{c},\mathbf{G}_{o})$ using the equations in (15).
Once $\mathbf{G}_{c}$ is known, the question if $\mathbf{G}$ can be recovered arises. Conditions under which this is possible are in the literature called generic network identifiability, and has been studied in e.g. [21]. It is straightforward to conclude that the closed loop transfer function from $\mathbf{r}$ to $\mathbf{x}_{o}$ is given by
$\mathbf{G}_{c}=J_{eo}(I-\mathbf{G}\Upsilon)^{-1}\mathbf{G}\Omega+J_{ro},$ (16)
which may be written as
$J_{eo}\operatorname{adj}(I-\mathbf{G}\Upsilon)\mathbf{G}\Omega+\left(J_{ro}-\mathbf{G}_{c}\right)\det(I-\mathbf{G}\Upsilon)=0.$ (17)
Mathematically, the question is if $\mathbf{G}$ can be solved for uniquely in this equation. For given values of $\mathbf{G}_{c}$, $\Upsilon$ and $\Omega$, this equation is a polynomial equation, and hence the answer to the question can be obtained by computing a Gröbner basis of the ideal generated by the elements of the matrix
$J_{eo}\operatorname{adj}(I-\mathbf{G}\Upsilon)\mathbf{G}\Omega+\left(J_{ro}-\mathbf{G}_{c}\right)\det(I-\mathbf{G}\Upsilon)$ (18)
viewed as polynomials in the entries of $\mathbf{G}$. If the reduced Gröbner basis is 1, then there is no solution $\mathbf{G}$ to the equation, and otherwise, there is, [7]. In the latter case, we can investigate generic network identifiability by computing all the solutions using any of the algorithms presented in [15]. If there is only one solution we have generic network identifiability. We summarize this result in the following theorem.
Theorem 8
There is a solution $\mathbf{G}$ to (17) for given values of $\mathbf{G}_{c}$, $\Upsilon$ and $\Omega$ if and only if the reduced Gröbner basis for (18) is not equal to one.
Remark 4
When there is a solution, the uniqueness is investigated by computing all solutions using any of the algorithms in e.g. [15].
Remark 5
Under Assumption A1, $(a,b)$ is uniquely given from knowledge of $\mathbf{G}$.
Consider the example of a network of dynamical systems as described by the block diagram in Figure 1. Assume that we only measure $\mathbf{u}^{3}$. Some calculations show that
$(1-\mathbf{G}^{1}\mathbf{G}^{3})\mathbf{u}^{3}$ $=\mathbf{G}^{1}\mathbf{r}^{1}+\mathbf{G}^{1}\mathbf{G}^{2}\mathbf{r}^{2}+\mathbf{r}^{3}$ (19)
$+\mathbf{H}^{1}\mathbf{e}^{1}+\mathbf{G}^{1}\mathbf{H}^{2}\mathbf{e}^{2}+\mathbf{G}^{1}\mathbf{H}^{3}\mathbf{e}^{3},$
and from this we realize that (17) may be written as
$\begin{bmatrix}\mathbf{G}^{1}&\mathbf{G}^{1}\mathbf{G}^{2}&1\end{bmatrix}-(1-\mathbf{G}^{1}\mathbf{G}^{3})\mathbf{G}_{c}=0.$
From this, we may derive that
$\mathbf{G}^{1}=\frac{\mathbf{G}_{c}^{1}}{\mathbf{G}_{c}^{3}},\quad\mathbf{G}^{2}=\frac{\mathbf{G}_{c}^{2}}{\mathbf{G}_{c}^{1}},\quad\mathbf{G}^{3}=\frac{\mathbf{G}_{c}^{3}-1}{\mathbf{G}_{c}^{1}}.$
Hence we have generic network identifiability for this example. Notice that we cannot recover $\mathbf{H}^{i}$ since we do not know $\mathbf{e}^{i}$ like we know $\mathbf{r}^{i}$. The condition for generic network identifiability is valid irrespective of if we use a direct or an indirect method for system identification. In an indirect method, we identify $\mathbf{G}_{c}$ and use the above equations to solve for $\mathbf{G}$. In a direct method we directly identify $\mathbf{G}$ since the transfer function $\mathbf{G}_{c}$ is parameterized in terms of the parameters of the transfer function $\mathbf{G}$.
$\mathbf{G}^{2}$$\mathbf{G}^{3}$$\mathbf{G}^{1}$$\mathbf{r}^{1}$$\mathbf{r}^{2}=u^{2}$$\mathbf{r}^{3}$$\mathbf{v}^{1}$$\mathbf{v}^{2}$$\mathbf{v}^{3}$$\mathbf{y}^{2}$$\mathbf{y}^{3}$$\mathbf{u}^{1}$$\mathbf{y}^{1}$$\mathbf{u}^{3}$
Figure 1: Block diagram for a dynamic network where $\mathbf{v}^{i}=\mathbf{H}^{i}\mathbf{e}^{i}$.
Assuming generic network identifiability, we can thus identify $\mathbf{G}$. However, we are not able to identify $\mathbf{H}$ in general. We cannot even recover $\mathbf{H}_{c}$ in general. It is not enough to assume that $\mathbf{H}_{c}$ should be minimum phase. This is obvious from the above example. Because of this we can only hope to prove consistency for estimates of $(a,b)$, and not for estimates of $(c,\lambda)$. However, for special cases we may sometimes prove more, c.f. Remark 8. Since generic network identifiability is a purely algebraic question that can be investigated separately, we introduce the following assumption.
  1. A5.
    The equation in (17) has a unique solution $\mathbf{G}$.
VI Maximum Likelihood Estimation
There are two possible ways of defining an ML problem for system identification. Already in [27] it is mentioned that one possible formulation is to use the innovation from in (4), and make use the fact that the innovations $\epsilon_{k}$ are independent random variables with zero mean and covariance $\Sigma_{\epsilon}$. Hence the pdf of the innovations over the time interval one to $N$ is just the product of the pdfs for each time instant, and by considering the negative logarithm of this function we obtain the optimization problem
minimize $\frac{1}{2}\sum_{k=1}^{N}\left(\epsilon_{k}^{T}\Sigma_{\epsilon}^{-1}\epsilon_{k}+\ln\det\Sigma_{\epsilon}\right)$ (20)
subject to $\hat{\xi}_{k+1}=F_{c}\hat{\xi}_{k}+G_{r}r_{k}+K\epsilon_{k}$
$x_{o,k}=H_{o}\hat{\xi}_{k}+J_{ro}r_{k}+\epsilon_{k}$
[…]
with variables $(\theta,\hat{\xi},\epsilon,\Sigma,K,\Sigma_{\epsilon})$, and where $\xi_{1}$ is a free variable. Notice that $\Sigma_{\epsilon}$ is invertible by Lemma 1 and that the algebraic Riccati equation in (3) is also a constraint in the above optimization problem. The innovations are functions of the observations $x_{o}=(x_{o,1},\ldots,x_{o,N})$ through the constraints. Since $\hat{\xi}_{1}$ is a free variable, it can be argued that this is not a true ML formulation.
An alternative formulation of the ML problem is based on the fact that the pdf for the observations can be factorized as
$p(x_{o})=\prod_{k=1}^{N}p_{k}(x_{o,k}|x_{o,k-1},\ldots,x_{o,1}).$
Let us introduce the time-varying Kalman filter
$\check{\xi}_{k+1}$ $=F_{c}\check{\xi}_{k}+G_{r}r_{k}+K_{k}(x_{o,k}-H_{o}\check{\xi}_{k}-J_{ro}r_{k})$
$\check{x}_{o,k}$ $=H_{o}\check{\xi}_{k}+J_{ro}r_{k}$
where $K_{k}$ is obtained from the recursion
$\begin{bmatrix}I&K_{k}\\ 0&I\end{bmatrix}\begin{bmatrix}\Sigma_{k+1}&0\\ 0&\Sigma_{\epsilon,k}\end{bmatrix}\begin{bmatrix}I&K_{k}\\ 0&I\end{bmatrix}^{T}$ (21)
$\qquad=\begin{bmatrix}F_{c}&G_{e}\\ H_{o}&J_{eo}\end{bmatrix}\begin{bmatrix}\Sigma_{k}&0\\ 0&\Sigma_{e}\end{bmatrix}\begin{bmatrix}F_{c}&G_{e}\\ H_{o}&J_{eo}\end{bmatrix}^{T},$
with $\check{\xi}_{1}$ and $\Sigma_{1}$ given, which are the expected value and covariance of $\xi_{1}$, respectively.
If we assume that $r_{k}$ has been zero for $k\leq 0$, then $\check{\xi}_{1}=0$ and $\Sigma_{1}=P$, where $P$ is the solution of the algebraic Lyapunov equation
$P=F_{c}PF_{c}^{T}+G_{e}\Sigma_{e}G_{e}^{T}.$ (22)
Associated with the time-varying Kalman-filter are the non-stationary innovations $\check{\epsilon}_{k}=x_{o,k}-\check{x}_{o,k}$, and the non-stationary innovation-form
$\check{\xi}_{k+1}$ $=F_{c}\check{\xi}_{k}+G_{r}r_{k}+K_{k}\check{\epsilon}_{k}$
$x_{o,k}$ $=H_{o}\check{\xi}_{k}+J_{ro}r_{k}+\check{\epsilon}_{k},$
with $\check{\xi}_{1}=0$, and where $\check{\epsilon}_{k}$ is a sequence of independent zero mean random variables with covariances $\Sigma_{\epsilon,k}$. It now holds that $p_{k}(x_{o,k}|x_{o,k-1},\ldots,x_{o,1})$ are pdfs for Gaussian random variables with mean $\check{x}_{o,k}$ and covariance $\Sigma_{\epsilon,k}$. By taking the logarithm of the factorized version of the pdf for the observations we see that the ML estimation problem is equivalent to the optimization problem
minimize $\frac{1}{2}\sum_{k=1}^{N}\left(\check{\epsilon}_{k}^{T}\Sigma_{\epsilon,k}^{-1}\check{\epsilon}_{k}+\ln\det\Sigma_{\epsilon,k}\right)$ (23)
subject to $\check{\xi}_{k+1}=F_{c}\check{\xi}_{k}+G_{r}r_{k}+K_{k}\check{\epsilon}_{k}$
$x_{o,k}=H_{o}\check{\xi}_{k}+J_{ro}r_{k}+\check{\epsilon}_{k}$
[…]
with variables $(\theta,\check{\xi},\check{\epsilon},P,\Sigma_{k},K_{k},\Sigma_{\epsilon,k})$, and where $\check{\xi}_{1}=0$. Notice that $\Sigma_{\epsilon,k}$ is invertible by Lemma 13. Both the algebraic Lyapunov equation in (22) as well as the Riccati recursion in (21) are constraints in the optimization problem.
Neither of the above formulations are very good formulations of the system identification problem for numerical optimization, since the constraints depend on the parameter $\theta$ via the solution to either an algebraic Lyapunov equation or an algebraic Riccati equation. However, we will see that they are good formulations of the problem to prove consistency of the estimates. For numerical optimization we present an equivalent formulation of the first problem formulation in a later section where the algebraic Riccati equation is not present as a constraint.
VII Consistency
Consistency of ML estimation for system identification has a long history. The first investigation was performed in [27], where single-input single-output ARMAX models were investigated. Later references are e.g. [16, 20] and the references therein. The first reference of these two references shows convergence with probability one of parameters for linear, possibly time-varying, systems—hence not necessarily ergodic. It considers general time-varying predictors. The second reference considers multiple-input multiple-output ARMAX models under ergodic assumptions. It does in detail investigate the time-varying predictor based on the conditional mean with respect to the observed data. We will investigate consistency for the time-varying predictor and obtain consistency for the time-invariant predictor for free. Our proof is a generalization of the results in [20]. Introduce the log-likelihood function
$L_{N}(\theta)=\frac{1}{2N}\sum_{k=1}^{N}\left(\check{\epsilon}_{k}^{T}\Sigma_{\epsilon,k}^{-1}\check{\epsilon}_{k}+\ln\det\Sigma_{\epsilon,k}\right)+\frac{n_{o}}{2}\ln(2\pi),$
where $\check{\epsilon}_{k}$ was defined in the previous section. Notice that both $\check{\epsilon}_{k}$ and $\Sigma_{\epsilon,k}$ are functions of $\theta$, i.e. we may write $\check{\epsilon}_{k}(\theta)=x_{o,k}-\check{x}_{o,k}(\theta)$ and $\Sigma_{\epsilon,k}(\theta)$. We emphasize that $x_{o,k}$ is obtained from a model with parameter value $\theta_{0}$ which is in general not equal to $\theta$. We denote with $\hat{\theta}_{N}$ the optimizing argument of $L_{N}(\theta)$. We want to show that $(\hat{a}_{N},\hat{b}_{N})\rightarrow(a_{0},b_{0})$ as $N\rightarrow\infty$ with probability one, i.e., we want to show what is called consistency.
The challenging part of proving consistency is to show that
$\sup_{\theta\in\Theta}\left|L_{N}(\theta)-L(\theta)\right|\rightarrow 0,\quad N\rightarrow\infty\;\mathrm{w.p.1}$
where $L(\theta)=-\operatorname{\mathbb{E}}\!\left[\,\ln f(\epsilon_{1}(\theta);\theta)\,\right]$, and where
$f(\epsilon;\theta)=\frac{1}{\sqrt{(2\pi)^{n_{0}}\det\Sigma_{\epsilon}(\theta)}}\exp\left(-\frac{1}{2}\epsilon^{T}\Sigma_{\epsilon}(\theta)^{-1}\epsilon\right).$
Notice that the time index 1 in $\epsilon_{1}(\theta)$ is arbitrary since $\epsilon_{k}(\theta)$ is a stationary stochastic process. This uniform convergence with probability one is proven in Lemma 16. For the time-invariant predictor formulation of the ML problem this result follows directly from Lemma 15. Denote by $\Theta_{\min}$ the set of $\theta\in\Theta$ that minimizes $L(\theta)$. It then follows, c.f. [17, Theorem 8.2], that
$\inf_{\theta\in\Theta_{\min}}\left\|\hat{\theta}_{N}-\theta\right\|_{2}\rightarrow 0,\quad N\rightarrow\infty\;\mathrm{w.p.\,\,1}.$
We want to show that the set $\Theta_{\min}$ contains $\theta_{0}$. Notice that
$L(\theta)=\frac{1}{2}\operatorname{tr}Q\Sigma_{\epsilon}(\theta)^{-1}+\frac{1}{2}\log\det\Sigma_{\epsilon}(\theta)+\text{constant},$
where $Q=\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta)\epsilon_{1}(\theta)^{T}\,\right]$. The function $L$ is a strictly convex function in $\Sigma_{\epsilon}(\theta)^{-1}$. The gradient of the function with respect to $\Sigma_{\epsilon}(\theta)^{-1}$ is $Q-\Sigma_{\epsilon}(\theta)$ from which we conclude that $\Sigma_{\epsilon}(\theta)=Q$ is the unique minimizer of $L(\theta)$ with minimal value
$\ln\det\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta)\epsilon_{1}(\theta)^{T}\,\right]+n_{o}+\text{constant}.$
From the fact that the mean square error is minimized by the Kalman filter, it holds that
$\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta)\epsilon_{1}(\theta)^{T}\,\right]\succeq\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta_{0})\epsilon_{1}(\theta_{0})^{T}\,\right]=\Sigma_{\epsilon}(\theta_{0})$ (24)
for all $\theta\in\Theta$. Therefore, we have
$L(\theta)$ $\geq\ln\det\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta)\epsilon_{1}(\theta)^{T}\,\right]+n_{o}+\text{constant}$
$\geq L(\theta_{0})=\ln\det\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta_{0})\epsilon_{1}(\theta_{0})^{T}\,\right]+n_{0}+\text{constant},$
where the second inequality follows from (24). Equality between $L(\theta)$ and $L(\theta_{0})$ holds only if
$\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta)\epsilon_{1}(\theta)^{T}\,\right]=\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta_{0})\epsilon_{1}(\theta_{0})^{T}\,\right],$
since if the determinant of two positive definite matrices are the same and they satisfy (24), then they are equal.
What can be said about $\theta$? It holds for any $\theta$ that
$\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta)\epsilon_{1}(\theta)^{T}\,\right]-\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta_{0})\epsilon_{1}(\theta_{0})^{T}\,\right]$
$\quad\quad\quad=\operatorname{\mathbb{E}}\!\left[\,\left(\epsilon_{1}(\theta)-\epsilon_{1}(\theta_{0})\right)\epsilon_{1}(\theta_{0})^{T}\,\right]$
$\quad\quad\quad\quad+\frac{1}{2}\operatorname{\mathbb{E}}\!\left[\,\left(\epsilon_{1}(\theta)-\epsilon_{1}(\theta_{0})\right)\left(\epsilon_{1}(\theta)-\epsilon_{1}(\theta_{0})\right)^{T}\,\right].$
Moreover $\epsilon_{1}(\theta)-\epsilon_{1}(\theta_{0})=\hat{x}_{o,1}(\theta_{0})-\hat{x}_{o,1}(\theta)$, where $\hat{x}_{o,1}(\theta_{0})$ and $\hat{x}_{o,1}(\theta)$ are independent of $\epsilon_{1}(\theta_{0})$. Hence by Lemma 6 we have for $\theta$ that minimizes $L(\theta)$ that $\mathbf{W}\left(e^{i\omega};\theta\right)=\mathbf{W}\left(e^{i\omega};\theta_{0}\right)$ for almost all $\omega$. We then obtain the following theorem from Lemma 7 and Remark 5.
Theorem 9
Assume that assumptions A1–A5 hold and that $\mathbf{C}^{i}(z)\neq 0$ for $|z|\geq 1$ and for all $i$ and uniformly in $\theta\in\Theta$. Then
[…]
i.e. $(a,b)$ are estimated consistently for either of the ML formulations.333 In practice we only need to assume that $\mathbf{C}^{i}(z)\neq 0$ for $|z|=1$, uniformly in $\theta\in\Theta$. What will happen is that we will identify $\mathbf{C}$ polynomials that have zeros mirrored in the unit circle if the polynomials that generated the data had zeros outside the unit circle. Another way to phrase it is that we cannot see in the observed data if a zero is inside our outside the unit circle, and hence we may just as well assume it to be inside.
Remark 6
From Theorem 9.1 and (9.31) in [17] it follows that $\sqrt{N}\left((\hat{a}_{N},\hat{b}_{N})-(a_{0},b_{0})\right)$ is asymptotically normal with zero mean, and with a covariance that equals the limit of the Cramér-Rao lower bound, i.e. the estimator is efficient. [^4] The proof of this is immediate by just considering the subvector $(a,b)$ of $\theta$ in the proofs in [17].
Remark 7
Notice that we are not assuming that $e_{k}$ is Gaussian in order to prove consistency. It is enough that it is a stationary stochastic process with zero mean and bounded second moment.
Remark 8
In case we instead of considering $\theta=(a,b,c,\lambda)$ as parameters consider $\bar{\theta}=(a,b,K,\Sigma_{\epsilon})$ to be the parameters we have that $\bar{\theta}$ sometimes can be estimated consistently. Notice that this does not mean that we use $\bar{\theta}$ as optimization variable.
Since the predictor is estimated consistently we can from the predictions and observations compute innovations, and from a realization of these we may estimate their covariance, which is $\Sigma_{\epsilon}$. This can always be done consistently. Let us now look at the special cases in Remark 2. When $T_{o}=\begin{bmatrix}I&0\end{bmatrix}$ it holds that $\Sigma_{\epsilon}=\Sigma_{e}$, and hence it is possible to estimate $\lambda$ consistently.
By Lemma 7 the observer form transfer function in (14) is estimated consistently. Since we know that $(a,b)$ are estimated consistently it only remains to investigate if $K$ is also estimated consistently. The Markov parameters for the transfer function are also estimated consistently and we denote them with $M_{k}$. They satisfy $H_{o}F_{c}^{k-1}K=M_{k}$ for $k\geq 1$. Hence we obtain the equation
$\begin{bmatrix}H_{o}\\ H_{o}F_{c}\\ \vdots\\ H_{o}F_{c}^{n-1}\end{bmatrix}K=\begin{bmatrix}M_{1}\\ M_{2}\\ \vdots\\ M_{n}\end{bmatrix}$
for $K$. If the observability matrix has full column rank, then we can solve for $K$, and then $K$ is also estimated consistently. This is however not true in general.
When the conditions of Theorem 2 are satisfied it holds that $K$ is the product of $G_{e}=C+B\Upsilon$ and an invertible matrix. Hence if both $b$ and $K$ are estimated consistently, then $c$ is also estimated consistently.
VIII Discussion
As already pointed out, the optimization problems for solving the ML problem are not very tractable, since they in general involve constraints either in terms of an algebraic Riccati equation or an algebraic Lyapunov equation. It is tempting to just remove the algebraic Riccati equation in (20) as a constraint. This will result in a relaxation of the optimization problem. This can potentially result in overfitting, since there are more free parameters to optimize than what there should be. There are however cases for when the Riccati equation is not needed. One case is discussed in Theorem 4. The assumptions of that theorem hold for the network in Figure 1 for the case when we observe $x_{o,k}=(u^{1}_{k},u^{3}_{k})$. The predictor transfer functions in (12) are given by
$\hat{\mathbf{u}}^{1}$ $=\mathbf{r}^{1}+\mathbf{G}^{2}\mathbf{r}^{2}+\mathbf{G}^{3}\mathbf{u}^{3}$ (25)
$\hat{\mathbf{u}}^{3}$ $=\mathbf{G}^{1}\mathbf{u}^{1}+\mathbf{r}^{3}.$
Since $\Sigma=0$, it holds that $\Sigma_{\epsilon}=R=\operatorname{diag}(\lambda_{2}+\lambda_{3},\lambda_{1})$, and the ML problem is therefore equivalent to minimizing
$\frac{1}{2}\sum_{k=1}^{N}\left(\frac{\left(u^{3}_{k}-\hat{u}^{3}_{k}\right)^{2}}{\lambda_{1}}+\frac{\left(u^{1}_{k}-\hat{u}^{1}_{k}\right)^{2}}{\lambda_{2}+\lambda_{3}}+\ln(\lambda_{1})+\ln(\lambda_{2}\!+\!\lambda_{3})\!\right).$
Since the terms in the objective function depend on a disjoint set of parameters, the minimization problem is separable. This results in a PEM where the two functions
$\frac{1}{2}\sum_{k=1}^{N}\left(u^{3}_{k}-\hat{u}^{3}_{k}\right)^{2},\quad\frac{1}{2}\sum_{k=1}^{N}\left(u^{1}_{k}-\hat{u}^{1}_{k}\right)^{2}$
are minimized, and this is what is suggested in e.g. [22, 9]. Notice that only the sum $\lambda_{2}+\lambda_{3}$ can be estimated consistently.
For ARMAX models, where $a\neq c$, we have for the network in Figure 1
$\mathbf{u}^{1}$ $=\mathbf{r}^{1}+\mathbf{G}^{2}\mathbf{r}^{2}+\mathbf{G}^{3}\mathbf{u}^{3}+\mathbf{H}^{2}\mathbf{e}^{2}+\mathbf{H}^{3}\mathbf{e}^{3}$ (26)
$\mathbf{u}^{3}$ $=\mathbf{G}^{1}\mathbf{u}^{1}+\mathbf{r}^{3}+\mathbf{H}^{1}\mathbf{e}^{1}.$
When $a=c$, then $\mathbf{H^{i}}=1$, and the predictor in (25) is trivially obtained by erasing the terms containing $\mathbf{e}^{i}$. In case $\mathbf{H^{i}}\neq 1$, one first needs to perform a spectral factorization of $\mathbf{H}^{2}\mathbf{e}^{2}+\mathbf{H}^{3}\mathbf{e}^{3}$ before a predictor can be determined, i.e. one writes $\mathbf{H}^{2}\mathbf{e}^{2}+\mathbf{H}^{3}\mathbf{e}^{3}=\mathbf{\bar{H}}\mathbf{\bar{\epsilon}}$, where $\mathbf{\bar{H}}$ is parameterized in terms of $(a,c,\lambda)$ via an algebraic Riccati equation. In case one wants to use a PEM approach one instead needs to parameterize $\mathbf{\bar{H}}$ independently of $(a,c,\lambda)$. This will potentially result in overfitting. Our ML approach will automatically make sure that the algebraic Riccati equation is taken into consideration, and thus avoids overfitting.
The two equations in (26) cannot be used to form two independent predictors if any of the $\mathbf{e}^{i}$-terms in one of the equations would also be present in the other equation. This is the case if we e.g. add $\mathbf{y}^{2}=\mathbf{G}^{2}\mathbf{r}^{2}+\mathbf{H}^{2}\mathbf{e}^{2}$ to $\mathbf{u}^{3}$, which corresponds to a modified network. Then the spectral factorization involves the two equations jointly. Moreover, the transfer function $\mathbf{G}^{2}$ is then present in two equations and should be parameterized jointly in the two equations. A PEM approach based on Multiple-Input-Single-Output (MISO) models do not accommodate this. Our ML approach seamlessly does this.
For the network and observed signals considered in this example we obtain a predictor for which each of the transfer functions $\mathbf{G}^{i}$ appear linearly in (26). The PEM approach is strongly dependent on this. However, it is not always possible to obtain this linearity. For the case when only $\mathbf{u}^{3}$ is measured, we need to substitute the equation for $\mathbf{u}^{1}$ resulting in products of transfer functions, c.f. (19). It is also easy to verify that (11) does not hold. It should be stressed that we have generic network identifiability for this case, and hence our proposed ML approach can deal with this situation, for which the PEM approach is not applicable.
Let us now consider the simple example of a single feedback loop around one single transfer function. We have the equations
$\mathbf{y}=\mathbf{G}\mathbf{u}+\mathbf{e},\quad\mathbf{u}=\mathbf{y}+\mathbf{r}.$
We observe $(\mathbf{y},\mathbf{u})$. We notice that there is no noise present in the second equation. This is an example of what in [26] is called the "rank-reduced noise" case. We have a state space description of the closed loop system which is
$\xi_{k+1}$ $=(F+BH)\xi_{k}+Br_{r}+(B+C)e_{k}$
$\begin{bmatrix}y_{k}\\ u_{k}\end{bmatrix}$ $=\begin{bmatrix}1\\ 1\end{bmatrix}(H\xi_{k}+e_{k})+\begin{bmatrix}0\\ 1\end{bmatrix}r_{k}.$
For this example, Assumption A2 is not fulfilled. It holds that $y_{k}$ and $u_{k}$ can be expressed in one another via the equation $u_{k}=y_{k}+r_{k}$. According to Remark 1, one of the observations can be removed. Hence, if we keep the observation of $\mathbf{y}$, we obtain the equation
$\mathbf{y}=\mathbf{G}(\mathbf{y}+\mathbf{r})+\mathbf{e}.$
Since we always have that the innovation covariance $\Sigma_{\epsilon}$ is positive definite by Lemma 1 when fulfilling Assumption 2, we can by Remark 1 always transform ourselves to a non rank-reduced noise case.
Regarding the optimization problem in (23), the algebraic Lyapunov equation is not easy to remove either. One possibility is to remove it and let $\Sigma_{1}$ be a large matrix, e.g. a large number times the identity matrix. The time-varying predictor will for large $k$ converge to the correct predictor, which agrees with the time-invariant predictor resulting from the algebraic Riccati equation. In this way we obtain an approximation to the ML problem. Notice that the Riccati recursion for this ML formulation is a simple constraint to deal with, since it can be solved recursively once $\Sigma_{1}$ is known.
IX Equivalent Optimization Problem
Neither of the suggested approaches to remove the complicating constraints above are satisfactory, since they will in general not solve the ML problem. We will now show how we can formulate an optimization problem that is equivalent to the problem in (20) without using a predictor and hence without the need for an algebraic Riccati equation. This was first presented in [8] and summarized here for completeness.
We will assume that $y_{k}^{i}$, $u_{k}^{i}$, and $e_{k}^{i}$ are zero for all $k\leq 0$. [^5] The generalization to nonzero initial values is considered in e.g. [11, Section 12.3]. To simplify notation, we define $u^{i}=(u_{1}^{i},\ldots,u_{N}^{i})$, $y^{i}=(y_{1}^{i},\ldots,y_{N}^{i})$, and $e^{i}=(e_{1}^{i},\ldots,e_{N}^{i})$. Moreover, we define $y=(y^{1},\ldots,y^{M})$, $u=(u^{1},u^{2},\ldots,u^{M})$, $e=(e^{1},e^{2},\ldots,e^{M})$, $r=(r^{1},r^{2},\ldots,r^{M})$, and lower-triangular Toeplitz matrices $T_{a^{i}}\in\mathbb{R}^{N\times N}$, $T_{b^{i}}\in\mathbb{R}^{N\times N}$ and $T_{c^{i}}\in\mathbb{R}^{N\times N}$ whose first columns are
$\begin{bmatrix}1\\ a^{i}\\ 0\end{bmatrix},\quad\begin{bmatrix}0\\ b^{i}\\ 0\end{bmatrix},\quad\begin{bmatrix}1\\ c^{i}\\ 0\end{bmatrix}$,
respectively, c.f., [25]. This allows us to express the ARMAX models in (1) as
$T_{a^{i}}y^{i}=T_{b^{i}}u^{i}+T_{c^{i}}e^{i}.$ (27)
We also define two block-diagonal matrices,
$T_{y}$ $=\operatorname{blkdiag}(T_{y^{1}},\ldots,T_{y^{M}}),$
$T_{u}$ $=\operatorname{blkdiag}(T_{u^{1}},\ldots,T_{u^{M}}),$
where $T_{y^{i}}=T_{c^{i}}^{-1}T_{a^{i}}$ and $T_{u^{i}}=T_{c^{i}}^{-1}T_{b^{i}}$. This allows us to express the dynamic network model as an instance of
$Ax+b=\begin{bmatrix}e\\ 0\end{bmatrix}$, (28)
with
$A$ $=\begin{bmatrix}A_{1}\\ A_{2}\end{bmatrix}=\begin{bmatrix}T_{y}&-T_{u}\\ -\Upsilon\otimes I&I\end{bmatrix}(P\otimes I)$ (29)
$x$ $=(P\otimes I)^{T}\begin{bmatrix}y\\ u\end{bmatrix},\quad b=\begin{bmatrix}b_{1}\\ b_{2}\end{bmatrix}=-\begin{bmatrix}0\\ \Omega\otimes I\end{bmatrix}r,$
where $A_{1},A_{2}\in\mathbb{R}^{MN\times 2MN}$, and where $P\in\mathbb{R}^{2M\times 2M}$ is a permutation matrix that is defined such that the observed parts of $y$ and $u$ correspond to the leading entries of $x$. The vector $b\in\mathbb{R}^{2MN}$ is partitioned conformable to $A$. Notice that the pdf for $x$ is singular.
The matrix $A$ is square, and using the fact that the matrices $T_{a^{i}}$ and $T_{c^{i}}$ are non-singular for all $i$, we see that $T_{y}$ is non-singular. Thus, $A$ is non-singular if the Schur complement $T_{y}-T_{u}(\Upsilon\otimes I)$ is full rank. This is true if the closed loop system is well-posed. Notice that $A_{1}$ depends on the unknown model parameters $\theta$, but $A_{2}$ and $b$ do not depend on $\theta$. We will see that this allows us to rewrite the pdf for $x$ as a nonsingular pdf and an algebraic constraint.
Consider (28), where $x$ is partitioned as $x=(x_{o},x_{m})\in\mathbb{R}^{n_{o}+n_{m}}$, corresponding to observed and missing data, respectively, $A(\theta)\in\mathbb{R}^{2MN\times(n_{o}+n_{m})}$ and $b\in\mathbb{R}^{2MN}$ are partitioned conformably as
$A(\theta)=\begin{bmatrix}A_{1}(\theta)\\ A_{2}\end{bmatrix}=\begin{bmatrix}A_{1o}(\theta)&A_{1m}(\theta)\\ A_{2o}&A_{2m}\end{bmatrix},\ b=\begin{bmatrix}b_{1}\\ b_{2}\end{bmatrix}$.
Given a Singular Value Decomposition (SVD) of $A_{2m}$, i.e.,
$A_{2m}=\begin{bmatrix}U_{1}&U_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{1}&V_{2}\end{bmatrix}^{T},$
we can rewrite (28) as
$\begin{bmatrix}A_{1o}&\bar{A}_{1m1}&\bar{A}_{1m2}\\ \bar{A}_{2o1}&\Sigma_{1}&0\\ \bar{A}_{2o2}&0&0\end{bmatrix}\begin{bmatrix}x_{o}\\ \bar{x}_{m1}\\ \bar{x}_{m2}\end{bmatrix}+\begin{bmatrix}b_{1}\\ \bar{b}_{21}\\ \bar{b}_{22}\end{bmatrix}=\begin{bmatrix}e\\ 0\\ 0\end{bmatrix}$
where
$\bar{A}_{2o}$ $=\begin{bmatrix}\bar{A}_{2o1}\\ \bar{A}_{2o2}\end{bmatrix}=\begin{bmatrix}U_{1}^{T}\\ U_{2}^{T}\end{bmatrix}A_{2o},$ (30)
$\bar{A}_{1m}$ $=\begin{bmatrix}\bar{A}_{1m1}&\bar{A}_{1m2}\end{bmatrix}=A_{1m}\begin{bmatrix}V_{1}&V_{2}\end{bmatrix}$ (31)
and
$\bar{x}_{m}$ $=\begin{bmatrix}\bar{x}_{m1}\\ \bar{x}_{m2}\end{bmatrix}=\begin{bmatrix}V_{1}^{T}\\ V_{2}^{T}\end{bmatrix}x_{m},$ $\bar{b}_{2}$ $=\begin{bmatrix}\bar{b}_{21}\\ \bar{b}_{22}\end{bmatrix}=\begin{bmatrix}U_{1}^{T}\\ U_{2}^{T}\end{bmatrix}b_{2}.$
Using $\Sigma_{1}$ as a pivot, we can rewrite the system as
$\begin{bmatrix}\bar{A}_{1o}&0&\bar{A}_{1m2}\\ \bar{A}_{2o1}&\Sigma_{1}&0\\ \bar{A}_{2o2}&0&0\end{bmatrix}\begin{bmatrix}x_{o}\\ \bar{x}_{m1}\\ \bar{x}_{m2}\end{bmatrix}+\begin{bmatrix}\bar{b}_{1}\\ \bar{b}_{21}\\ \bar{b}_{22}\end{bmatrix}=\begin{bmatrix}e\\ 0\\ 0\end{bmatrix}$, (32)
where $\bar{A}_{1o}=A_{1o}-\bar{A}_{1m1}\Sigma_{1}^{-1}\bar{A}_{2o1}$ and $\bar{b}_{1}=b_{1}-\bar{A}_{1m1}\Sigma_{1}^{-1}\bar{b}_{21}$. The fact that $A$ is invertible implies that $\bar{A}_{2o2}$ has full row-rank, and hence there exists an orthogonal matrix $W=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}$ such that $\bar{A}_{2o2}W=\begin{bmatrix}\tilde{A}_{2o21}&0\end{bmatrix}$ with $\tilde{A}_{2o21}=\bar{A}_{2o2}W_{1}$ nonsingular. The matrix $W$ can be obtained by means of an SVD or an LQ decomposition of $\bar{A}_{2o2}$. We then define
$\begin{bmatrix}\tilde{A}_{1o}\\ \tilde{A}_{2o1}\end{bmatrix}$ $=\begin{bmatrix}\tilde{A}_{1o1}&\tilde{A}_{1o2}\\ \tilde{A}_{2o11}&\tilde{A}_{2o12}\end{bmatrix}=\begin{bmatrix}\bar{A}_{1o}\\ \bar{A}_{2o1}\end{bmatrix}W,$ (33)
$\bar{x}_{o}$ $=\begin{bmatrix}\bar{x}_{o1}\\ \bar{x}_{o2}\end{bmatrix}=W^{T}x_{o},$
which allows us to rewrite (32) as
$\begin{bmatrix}\tilde{A}_{1o1}&\tilde{A}_{1o2}&0&\bar{A}_{1m2}\\ \tilde{A}_{2o11}&\tilde{A}_{2o12}&\Sigma_{1}&0\\ \tilde{A}_{2o21}&0&0&0\end{bmatrix}\begin{bmatrix}\bar{x}_{o1}\\ \bar{x}_{o2}\\ \bar{x}_{m1}\\ \bar{x}_{m2}\end{bmatrix}+\begin{bmatrix}\bar{b}_{1}\\ \bar{b}_{21}\\ \bar{b}_{22}\end{bmatrix}=\begin{bmatrix}e\\ 0\\ 0\end{bmatrix}$.
Using $\tilde{A}_{2o21}$ as a pivot, we obtain the system
$\begin{bmatrix}0&\tilde{A}_{1o2}&0&\bar{A}_{1m2}\\ \tilde{A}_{2o11}&\tilde{A}_{2o12}&\Sigma_{1}&0\\ \tilde{A}_{2o21}&0&0&0\end{bmatrix}\begin{bmatrix}\bar{x}_{o1}\\ \bar{x}_{o2}\\ \bar{x}_{m1}\\ \bar{x}_{m2}\end{bmatrix}+\begin{bmatrix}\tilde{b}_{1}\\ \bar{b}_{21}\\ \bar{b}_{22}\end{bmatrix}=\begin{bmatrix}e\\ 0\\ 0\end{bmatrix}$,
where $\tilde{b}_{1}=\bar{b}_{1}-\tilde{A}_{1o1}\tilde{A}_{2o21}^{-1}\bar{b}_{22}$. Finally, we rewrite this system as two systems,
$\begin{bmatrix}\tilde{A}_{1o2}&\bar{A}_{1m2}\end{bmatrix}\begin{bmatrix}\bar{x}_{o2}\\ \bar{x}_{m2}\end{bmatrix}+\tilde{b}_{1}=e,$ (34)
$\begin{bmatrix}\tilde{A}_{2o11}&\Sigma_{1}\\ \tilde{A}_{2o21}&0\end{bmatrix}\begin{bmatrix}\bar{x}_{o1}\\ \bar{x}_{m1}\end{bmatrix}+\begin{bmatrix}\bar{b}_{21}+\tilde{A}_{2o12}\bar{x}_{o2}\\ \bar{b}_{22}\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}$. (35)
From this we realize that only $\bar{x}_{o2}$ and $\bar{x}_{m2}$ are directly related to $e$. The matrix $\begin{bmatrix}\tilde{A}_{1o2}&\bar{A}_{1m2}\end{bmatrix}$ has full row-rank, so if it is a square matrix, then it is also invertible, and otherwise we can make use of column compression to further reduce the number of variables. The equation in (34) describes a nonsingular pdf for $(\bar{x}_{o2},\bar{x}_{m2})$, and it is of the form for which it is possible to solve the ML problem based on only measuring $\bar{x}_{o2}$ as described in [13]. [^6] In this reference all $\lambda^{i}$ are equal, but it is a straightforward extension to consider the case when they are not equal. The fact that the observations in $\bar{x}_{o1}$ are not needed is related to the observations not needed in Remark 1.
X Numerical Results
Our implementation is based on the Python library JAX [3], which makes use of automatic differentiation to compute the partial derivatives of the cost function. The optimization problem associated with the ML approach is non-convex. To ensure convergence to a useful solution, an appropriate initialization strategy is therefore required. We first solve an ML problem where we consider $\mathbf{C}^{i}(z)=1$, i.e. we approximate each system with an ARX model and we also constrain all $\lambda^{i}$ to be the same. Then, we again solve an ML problem, but this time for an ARMAX model. We still constrain all $\lambda^{i}$ to be the same, and we initialize the solver with the value of $(a,b)$ we obtained in the previous step. This will provide us with initial values $(a,b,c,\lambda)$ for a third final step, where we do not constrain $\lambda^{i}$. We solve all the optimization problems with a trust-region method implemented via SciPy’s optimize module, see [24].
We will now illustrate some properties of the proposed method using a numerical example based on the network in Fig. 1. We generate 100 random models $\mathbf{M}(z;\theta)$ using the Matlab command drss. For each random model there are three sub-models with two inputs, one for $\mathbf{u}^{i}$ and one for $\mathbf{e}^{i}$, and one output for $\mathbf{y}^{i}$. We modify the transfer functions $\mathbf{H}^{i}(z;\theta)$ such that they have static gain equal to one. The order of all transfer functions is two. Models that are unstable are discarded. Similarly, any model for which the closed loop system $\mathbf{M}_{c}(z;\theta)$ has poles which are further away from the origin than 0.9 is also discarded. We generate different values for $\lambda_{i}$ for each model that are drawn from a uniform distribution on the interval $0.1\bar{\lambda},\bar{\lambda}]$, where $\bar{\lambda}=0.1$. For each model we draw $e_{k}^{i}$ from zero mean Gaussian distributions with variance $\lambda_{i}$ for $1\leq k\leq N$. The value $N$ is different for different experiments. The reference values $r_{k}^{i}$ are equal to $\pm 1$ with equal probability. All generated signals are independent of one another and in time. The average SNR is approximately given by $1/(0.55\bar{\lambda})\approx 25{\rm dB}$. We then generate observed data using
${\mathbf{x}}_{o}=\mathbf{G}_{c}(\theta_{0})\mathbf{r}+\mathbf{H}_{c}(\theta_{0})\mathbf{e},$ (36)
where we with subindex zero emphasize that this is a model from which we collect data. We use zero initial values. We will consider the cases when either $(\mathbf{u}^{1},\mathbf{u}^{3})$ or $\mathbf{u}^{3}$ are observed.
To validate the estimated models $\hat{\theta}$, the covariance
$\text{Cov}\left[\hat{\theta}\right$] $=\operatorname{\mathbb{E}}\!\left[\,(\hat{\theta}-\operatorname{\mathbb{E}}\hat{\theta})(\hat{\theta}-\operatorname{\mathbb{E}}\hat{\theta})^{T}\,\right],$
and the bias,
$\text{bias}\left[\hat{\theta}\right]=\operatorname{\mathbb{E}}\!\left[\,\hat{\theta}-\theta_{0}\,\right],$
of the estimates are computed numerically using averages, by performing 100 estimates considering different realizations of $e$ with the same values for $r$. We also compute the Mean Squared Error (MSE) given by
[…]
$=\operatorname{tr}\left(\text{Cov}\left[\hat{\theta}\right]\right)+\left\|\text{bias}\left[\hat{\theta}\right]\right\|_{2}^{2}.$
For computing fit, we generate new data according to (36), where the noise and reference values are different from the ones we used to obtain the model estimate. The simulation data for a model $\theta$ is defined as
$\hat{\mathbf{x}}_{o,s}=\mathbf{G}_{c}(\theta)\mathbf{r}.$
Notice that no noise realization is needed to generate the simulation data. The same reference value is used as the one used for generating the validation data. The prediction data for the model $\theta$ is defined as
$\hat{\mathbf{x}}_{o,p}=\mathbf{W}_{r}(\theta)\mathbf{r}+\mathbf{W}_{o}(\theta){\mathbf{x}}_{o}.$
Here there will be noise present, since the latter term contains the noise from the validation data. The fit is defined as
${\rm fit}(\hat{x}_{o})=1-\frac{\|\hat{x}_{o}^{i}-x_{o}^{i}\|_{2}}{\|x_{o}^{i}-\mathbf{1}\bar{x}_{o}^{i}\|_{2}},$
where $\bar{x}_{o}^{i}$ is the mean of $x_{o}^{i}$, and where superscript $i$ refers to the $i$th observed output. This definition can be used for either predicted or simulated data.
For the case when $(\mathbf{u}^{1},\mathbf{u}^{3})$ is observed, we can compare the ML method with the PEM method, for which we make use of the MISO models in (26), i.e.
$\mathbf{u}^{1}-\mathbf{r}^{1}$ $=\mathbf{G}^{3}\mathbf{u}^{3}+\mathbf{G}^{2}\mathbf{r}^{2}+\bar{\mathbf{H}}\bar{\mathbf{e}}$ (37)
$\mathbf{u}^{3}-\mathbf{r}^{3}$ $=\mathbf{G}^{1}\mathbf{u}^{1}+\mathbf{H}^{1}\mathbf{e}^{1}.$ (38)
To estimate the parameters of the model with PEM we use the System Identification Toolbox, [18], in MATLAB. We use the function armax to identify $(\mathbf{G}^{1},\mathbf{H}^{1})$ from (38) and the function bj to identify $(\mathbf{G}^{2},\mathbf{G}^{3},\bar{\mathbf{H}})$ from (37). We consider $\bar{\mathbf{H}}$ with orders one to four, and we choose the order that obtains the best fit value for prediction on the validation data. Notice that order four matches the order obtained from the spectral factorization of $\mathbf{H}^{2}\mathbf{e}^{2}+\mathbf{H}^{3}\mathbf{e}^{3}$, c.f. Section VIII. We cannot formulate $\mathbf{W}_{o}(\theta)$ for a model obtained from the PEM approach. This is because we do not obtain estimates of $(c,\lambda)$. Instead we use the MISO predictors obtained from (37) and (38),
$\hat{\mathbf{u}}_{1}$ $=\bar{\mathbf{H}}^{-1}(\mathbf{r}^{1}+\mathbf{G}^{3}\mathbf{u}^{3}+\mathbf{G}^{2}\mathbf{r}^{2})+(1-\bar{\mathbf{H}}^{-1})\mathbf{u}_{1}$
$\hat{\mathbf{u}}_{3}$ $=\mathbf{H}_{1}^{-1}(\mathbf{r}^{3}+\mathbf{G}^{1}\mathbf{u}^{1})+(1-\mathbf{H}_{1}^{-1})\mathbf{u}_{3}.$
For both PEM and the proposed MLE we set the tolerance in the optimization algorithms to be $10^{-5}$. Table I summarizes the results for 100 systems. We compare PEM with the proposed ML method, reporting (i) the number of cases that converged to a solution with positive fit values, (ii) the mean $\pm$ standard deviation of the fit values, calculated for the cases that converged, and (iii) the average time to convergence. Notice that it is not possible to directly identify any transfer functions with PEM when measuring only $u^{3}$, so the table includes only the results for the ML method for that case. We observe that the fits are about the same for both methods, but that PEM is able to obtain the solution faster. The order of $\bar{\mathbf{H}}$ for $N=500$ is 1, 2, 3, and 4 for 31, 39, 20, and 10 of the 100 models, respectively. For $N=50$ the numbers are 58, 19, 10, and 13, respectively. This shows that PEM especially for few data needs to choose a low order of $\bar{\mathbf{H}}$ in order not to overfit. Table II presents the results for fit, covariance, bias and MSE for the first of the 100 randomly generated systems. The fit values are computed from the model obtained from the average of $\hat{\theta}$ over the 100 noise realizations. Regarding fit, the PEM and the ML method have the same performance. Regarding MSE, the ML method has slightly smaller values as compared to the PEM. The only case when the PEM outperforms the ML method is with respect to covariance for the case of $N=50$. This indicates that PEM provides a biased estimate for few data, probably due to overfitting of the noise model.
There is a small degradation in fit with respect to predicted data when only one signal is measured as compared to the case when two signals are measured. There is also a degradation with respect MSE, covariance and bias of the estimates. However, the overall conclusion is that the ML approach is able to successfully obtain a good model by only measuring $\hat{u}^{3}$. [^7] Our code and data are available on GitHub at: https://github.com/Jvictormata/mle_dyn_net
TABLE I: Comparison of the PEM and ML for 100 random systems: mean fit values for both simulation and prediction data.
N Observed signals PEM ML
$100\times(\text{avg. Fit}\pm\text{std})$ conv. avg. time (s) $100\times(\text{avg. Fit}\pm\text{std})$ conv. avg. time (s)
$\hat{u}_{1_{\text{pred}}}$ $\hat{u}_{1_{\text{sim}}}$ $\hat{u}_{3_{\text{pred}}}$ $\hat{u}_{3_{\text{sim}}}$ $\hat{u}_{1_{\text{pred}}}$ $\hat{u}_{1_{\text{sim}}}$ $\hat{u}_{3_{\text{pred}}}$ $\hat{u}_{3_{\text{sim}}}$
50 $u^{3}$ 77.34 $\pm$ 13.05 69.32 $\pm$ 18.04 94 15.75
500 81.90 $\pm$ 13.92 75.13 $\pm$ 15.45 92 122.83
50 $u^{1}$ and $u^{3}$ 84.36 $\pm$ 11.46 74.58 $\pm$ 14.97 84.86 $\pm$ 9.96 69.49 $\pm$ 42.48 100 1.18 84.89 $\pm$ 10.19 74.88 $\pm$ 15.26 86.21 $\pm$ 8.48 70.64 $\pm$ 33.60 100 15.88
500 87.57 $\pm$ 8.12 76.56 $\pm$ 15.52 87.85 $\pm$ 7.52 74.11 $\pm$ 17.35 100 0.64 87.09 $\pm$ 8.15 76.92 $\pm$ 15.40 87.94 $\pm$ 7.40 74.91 $\pm$ 16.46 94 132.06
TABLE II: Comparison of the PEM and ML for system 1: fit values for both simulation and prediction data, and trace and max. eigenvalue of the covariance matrix.
N Observed signals PEM ML
$100\times\text{Fit}$ Cov. matrix $(a,b)$ $\|\text{bias}(a,b)\|_{2}^{2}$ MSE(a,b)(a,b) $100\times\text{Fit}$ Cov. matrix $(a,b)$ $\|\text{bias}(a,b)\|_{2}^{2}$ MSE(a,b)(a,b)
$\hat{u}_{1_{\text{pred}}}$ $\hat{u}_{1_{\text{sim}}}$ $\hat{u}_{3_{\text{pred}}}$ $\hat{u}_{3_{\text{sim}}}$ Trace Max. eigval. $\hat{u}_{1_{\text{pred}}}$ $\hat{u}_{1_{\text{sim}}}$ $\hat{u}_{3_{\text{pred}}}$ $\hat{u}_{3_{\text{sim}}}$ Trace Max. eigval.
50 $u^{3}$ 85.66 81.77 7.42e-01 4.44e-01 2.79e-01 1.02
500 88.05 84.15 3.77e-01 3.59e-01 2.46e-01 6.23e-01
50 $u^{1}$ and $u^{3}$ 87.37 82.36 87.14 81.48 5.11e-01 4.65e-01 2.96e-01 8.08e-01 87.06 82.25 87.13 81.46 5.69e-01 5.41e-01 1.09e-02 5.80e-01
500 88.12 81.00 89.16 84.18 5.84e-01 5.82e-01 2.32e-03 5.87e-01 88.03 81.00 89.16 84.18 4.14e-01 4.13e-01 1.75e-02 4.32e-01
XI Conclusions
In this paper we have investigated ML estimation methods for identification of dynamical systems in networks. We have considered both time-invariant and time-varying predictors, and we have proven consistency of the ML methods. They are also efficient. Moreover, we have formulated the resulting optimization problem as an unconstrained problem that can be solved efficiently. In addition to this we have derived necessary and sufficient conditions for network identifiability in terms of Gröbner basis. We see in our numerical investigations that the ML approach matches the PEM when both methods can be applied and, critically, remains effective for cases where the PEM is unable to identify any transfer functions. The ML method is applicable whenever we have network identifiability. When the PEM method can be used it will save computational time. However, notice that it might especially for short data sequences result in overfitting and biased estimates.
XII Appendix
XII-A Uniform Convergence of Riccati Recursion
Lemma 10
Assume that A2 holds and that $\mathbf{C}^{i}(z)$ are not equal to zero for $|z|\geq 1$. Then it holds that $(\bar{F}_{c},H_{o})$ is detectable and $(\bar{F}_{c},\bar{Q}^{1/2})$ is stabilizable.
Proof:
Since $\theta\in\Theta$ it holds that $F_{c}$ has all eigenvalues strictly inside the unit circle. Hence $(F_{c},H_{o})$ is trivially detectable and consequently $(\bar{F}_{c},H_{o})$ is also detectable, since detectability is invariant under output injection.
Let $\bar{G}_{e}=G_{e}\Sigma_{e}^{1/2}$ and $\bar{J}_{eo}=J_{eo}\Sigma^{1/2}_{e}$. Consider
$\bar{\mathbf{P}}(z)=\mathbf{P}(z)\begin{bmatrix}I&0\\ 0&\Sigma^{1/2}_{e}\end{bmatrix}=\begin{bmatrix}-zI+F_{c}&\bar{G}_{e}\\ H_{o}&\bar{J}_{eo}\end{bmatrix}$.
Since $\mathbf{C}^{i}(z)$ are not equal to zero for $|z|\geq 1$, it follows similarly as in the proof of Lemma 1 that $\bar{\mathbf{P}}(z)$ has full row rank for $|z|\geq 1$. Let $U$ be an orthonormal matrix such that
$\bar{\mathbf{P}}(z)U=\begin{bmatrix}-zI+F_{c}&\tilde{G}_{e1}&\tilde{G}_{e2}\\ H_{o}&\tilde{J}_{eo}&0\end{bmatrix}$
where $\tilde{J}_{eo}$ is invertible. Now we have that
$\begin{bmatrix}I&-\tilde{G}_{e1}\tilde{J}_{eo}^{-1}\\ 0&I\end{bmatrix}\!\bar{\mathbf{P}}(z)U\!=\!\begin{bmatrix}-zI\!+F_{c}-\tilde{G}_{e1}\tilde{J}_{eo}^{-1}H_{o}&0&\tilde{G}_{e2}\\ H_{o}&\tilde{J}_{eo}&0\end{bmatrix}$.
From this it follows by the PBH test for stabilizability that $(F_{c}-\tilde{G}_{e1}\tilde{J}_{eo}^{-1}H_{o},\tilde{G}_{e2})$ is stabilizable. We have $S=\bar{G}_{e}\bar{J}_{eo}^{T}=\tilde{G}_{e1}\tilde{J}_{eo}^{T}$ and $R=\bar{J}_{eo}\bar{J}_{eo}^{T}=\tilde{J}_{eo}\tilde{J}_{eo}^{T}$ and hence $\bar{F}_{c}=F_{c}-SR^{-1}H_{o}=F_{c}-\tilde{G}_{e1}\tilde{J}_{eo}^{-1}H_{o}$. Moreover, $Q=\bar{G}_{e}\bar{G}_{e}^{T}=\tilde{G}_{e1}\tilde{G}_{e1}^{T}+\tilde{G}_{e2}\tilde{G}_{e2}^{T}$ and hence it holds that $\bar{Q}=Q-SR^{-1}S^{T}=\tilde{G}_{e2}\tilde{G}_{e2}^{T}$. From this it follows that $(\bar{F}_{c},\bar{Q}^{1/2})$ is stabilizable. ∎
Multiply (7) with $\begin{bmatrix}I&-K\\ 0&I\end{bmatrix}$ from the left and its transpose from the right to obtain the equivalent equations
$\Sigma$ $=(F_{c}-KH_{o})\Sigma(F_{c}-KH_{o})^{T}$
$\quad+Q-KS^{T}-SK^{T}+KRK^{T}$
$0$ $=(F_{c}-KH_{o})\Sigma H_{o}^{T}+S-KR$
$\Sigma_{\epsilon}$ $=H_{o}\Sigma H_{o}^{T}+R,$
which equivalently may be written as
$\Sigma$ $=(F_{c}-KH_{o})\Sigma(F_{c}-KH_{o})^{T}+Q-SR^{-1}S^{T}$ (39)
$\quad+\left(K-SR^{-1}\right)R\left(K-SR^{-1}\right)$
$K$ $=(F_{c}\Sigma H_{o}^{T}+S)(H_{o}\Sigma H_{o}^{T}+R)^{-1}$
$\Sigma_{\epsilon}$ $=H_{o}\Sigma H_{o}^{T}+R.$
Define the functions
$\mathcal{R}(\Sigma,K)$ $=(F_{c}-KH_{o})\Sigma(F_{c}-KH_{o})^{T}+Q-SR^{-1}S^{T}$
$\quad+(K-SR^{-1})R(K-SR^{-1})^{T}$
$\mathcal{G}(\Sigma)$ $=(F_{c}\Sigma H_{o}^{T}+S)(H_{o}\Sigma H_{o}^{T}+R)^{-1}.$
Then, from (39), the Riccati recursion in (21) is equivalent to
$\Sigma_{k+1}=\mathcal{R}(\Sigma_{k},K_{k}),\quad K_{k}=\mathcal{G}(\Sigma_{k})$
and the algebraic Riccati equation in (3) is equivalent to
$\Sigma=\mathcal{R}(\Sigma,K),\quad K=\mathcal{G}(\Sigma).$
We have the following lemmas.
Lemma 11
Let $0\preceq\Sigma_{1}\preceq\bar{\Sigma}_{1}$, let $\bar{\Sigma}_{k+1}=\mathcal{R}(\bar{\Sigma}_{k},\bar{K}_{k})$ with $\bar{K}_{k}$ arbitrary and let $\Sigma_{k+1}=\mathcal{R}(\Sigma_{k},K_{k})$ with $K_{k}=\mathcal{G}(\Sigma_{k})$, then
$\Sigma_{k+1}\preceq\bar{\Sigma}_{k+1},\quad k=1,2,\dots.$
Proof:
See Theorem 2.1 in [5]. ∎
Lemma 12
Assume that A2-A3 hold uniformly in $\theta\in\Theta$. Then there exist constants $K_{\Sigma}>0$, finite, and $0<\lambda_{\Sigma}<1$ such that
$\sup_{\theta\in\Theta}\|\Sigma_{k}(\theta)-\Sigma(\theta)\|_{2}<K_{\Sigma}\lambda_{\Sigma}^{k},\quad k\geq 0$
where $\Sigma_{k}(\theta)$ is defined by the Riccati recursion in (21) with arbitrary initial condition $\Sigma_{1}(\theta)\succeq 0$, and where $\Sigma(\theta)$ is the solution of the algebraic Riccati equation in (3).
Proof:
Let $\bar{\Sigma}_{k+1}=\mathcal{R}(\bar{\Sigma}_{k},\bar{K}_{k})$ with $\bar{K}_{k}=\mathcal{G}(\Sigma)$ and $\Sigma$ being the solution of the algebraic Riccati equation, and let $\Sigma_{k+1}=\mathcal{R}(\Sigma_{k},K_{k})$ with $K_{k}=\mathcal{G}(\Sigma_{k})$. With $\bar{\Sigma}_{1}=\Sigma_{1}\succeq 0$, we obtain from Lemma 11 that
$\Sigma_{k+1}$ $\preceq(F_{c}-KH_{o})\bar{\Sigma}_{k}(F_{c}-KH_{o})^{T}+Q-SR^{-1}S^{T}$
$\quad+(K-SR^{-1})R(K-SR^{-1})^{-1},$
Subtracting $\Sigma=\mathcal{R}(\Sigma,K)$ and using (39) leads to
$\Sigma_{k+1}-\Sigma\preceq(F_{c}-KH_{o})(\bar{\Sigma}_{k}-\Sigma)(F_{c}-KH_{o})^{T}.$
Recursively applying this inequality leads to
$\Sigma_{k+1}-\Sigma\preceq(F_{c}-KH_{o})^{k}(\bar{\Sigma}_{1}-\Sigma)(F_{c}-KH_{o})^{Tk}.$
By Lemma 1 $F_{c}-KH_{o}$ has all its eigenvalues strictly inside the unit disc, and then, there exists $\lambda_{\Sigma}$ such that
$\|(F_{c}-KH_{o})\|_{2}^{2}\leq\lambda_{\Sigma}<1,$
By defining $K_{\Sigma}=\sup_{\theta\in\Theta}\|\bar{\Sigma}_{1}-\Sigma\|_{2}$, we have
$\|\Sigma_{k+1}-\Sigma\|_{2}\leq K_{\Sigma}\lambda_{\Sigma}^{k}.$
Because the same argument holds for all $\theta\in\Theta$, we have
$\sup_{\theta\in\Theta}\|\Sigma_{k}(\theta)-\Sigma(\theta)\|_{2}<K_{\Sigma}\lambda_{\Sigma}^{k}.$
Remark 9
Notice that the constants are independent of $\theta$, and hence this shows uniform geometric convergence.
Lemma 13
Assume that A2-A3 hold uniformly in $\theta\in\Theta$. Let $P(\theta)$ be the solution of the algebraic Lyapunov equation in (22), let $\Sigma_{k}(\theta)$ be the solution of the Riccati recursion in (21) with $\Sigma_{1}(\theta)=P(\theta)$, and let $\Sigma(\theta)$ be the solution of the algebraic Riccati equation in (3). Then it holds that $\Sigma_{k}(\theta)\succeq\Sigma(\theta)$ and that $\Sigma_{\epsilon,k}(\theta)\succeq\Sigma_{\epsilon}(\theta)\succ 0$ uniformly for all $k\geq 1$ and $\theta\in\Theta$.
Proof:
From the (1,1)-block in (3) we have
$\Sigma=F_{c}\Sigma F_{c}^{T}+G_{e}\Sigma_{e}G_{e}^{T}-K\Sigma_{\epsilon}K^{T}.$
and from $\Sigma_{1}=P$ we have by (22)
$\Sigma_{1}=F_{c}\Sigma_{1}F_{c}^{T}+G_{e}\Sigma_{e}G_{e}^{T}.$
Hence, by defining $\hat{P}=\Sigma_{1}-\Sigma$, we have
$\hat{P}=F_{c}\hat{P}F_{c}^{T}+K\Sigma_{\epsilon}K^{T}.$
By Lemma 1 it holds that $F_{c}$ has all eigenvalues strictly inside the unit circle. Since $K\Sigma_{\epsilon}K^{T}\succeq 0$, we have $\hat{P}\succeq 0$, leading to $\Sigma_{1}\succeq\Sigma$. From the equivalent formulation of the algebraic Riccati equation in (9) it follows by Lemma 1 in [2] that $\Sigma_{k}\succeq\Sigma$ for all $k\geq 1$. From the definitions of $\Sigma_{\epsilon}$ and $\Sigma_{\epsilon,k}$ it follows that $\Sigma_{\epsilon,k}\succeq\Sigma_{\epsilon}$ for all $k\geq 1$. From Lemma 1 we have that $\Sigma_{\epsilon}\succ 0$. All these results hold uniformly in $\theta\in\Theta$. ∎
Corollary 13.1
Assume that A2-A3 hold uniformly in $\theta\in\Theta$. Define $Q_{k}(\theta)=\Sigma_{\epsilon,k}(\theta)^{-1}-\Sigma_{\epsilon}(\theta)^{-1}$, $\Delta\Sigma_{\epsilon,k}(\theta)=\Sigma_{\epsilon,k}(\theta)-\Sigma_{\epsilon}(\theta)$ and $\Delta K_{k}(\theta)=K(\theta)-K_{k}(\theta)$. As a direct consequence of Lemmas 12 and 13, it holds:
  1. 1.
    $\|Q_{k}(\theta)\|_{2}\rightarrow 0$, geometrically as $k\rightarrow\infty$ and uniformly in $\theta\in\Theta$,
  2. 2.
    $\|\Sigma_{\epsilon}(\theta)^{-1}\Delta\Sigma_{\epsilon,k}(\theta)\|_{2}\rightarrow 0$, geometrically as $k\rightarrow\infty$ and uniformly in $\theta\in\Theta$,
  3. 3.
    $\|\Delta K_{k}(\theta)\|_{2}\leq M_{\Sigma}\lambda_{\Sigma}^{k}$,
with finite $M_{\Sigma}>0$, and $0<\lambda_{\Sigma}<1$.
Proof:
From Lemma 12 and recalling that
$\Sigma_{\epsilon,k}$ $=H_{o}\Sigma_{k}H_{o}^{T}+J_{eo}\Sigma_{e}J_{eo}^{T}$
$\Sigma_{\epsilon}$ $=H_{o}\Sigma H_{o}^{T}+\;J_{eo}\Sigma_{e}J_{eo}^{T},$
we can write
$\|\Sigma_{\epsilon,k}-\Sigma_{\epsilon}\|_{2}\leq\|H_{o}\|^{2}_{2}\|\Sigma_{k}-\Sigma\|_{2}\leq\|H_{o}\|^{2}_{2}K_{\Sigma}\lambda_{\Sigma}^{k}.$
Using the identity $A^{-1}-B^{-1}=A^{-1}(B-A)B^{-1}$, we can write
$\|Q_{k}\|_{2}=\|\Sigma_{\epsilon,k}^{-1}-\Sigma_{\epsilon}^{-1}\|_{2}\leq\|\Sigma_{\epsilon,k}^{-1}\|_{2}\,\|\Sigma_{\epsilon}^{-1}\|_{2}\,\|\Sigma_{\epsilon,k}-\Sigma_{\epsilon}\|_{2}.$
Since, from Lemma 13, $\|\Sigma_{\epsilon,k}(\theta)^{-1}\|_{2}$ and $\|\Sigma_{\epsilon}(\theta)^{-1}\|_{2}$ are both uniformly bounded for $\theta\in\Theta$ and $k\geq 0$, we have
$\|Q_{k}\|_{2}\leq\Bigl{(}\sup_{\begin{subarray}{c}k,\theta\end{subarray}}\|\Sigma_{\epsilon,k}^{-1}\|_{2}\,\|\Sigma_{\epsilon}^{-1}\|_{2}\,\|H_{o}\|^{2}_{2}\Bigr{)}\,K_{\Sigma}\lambda_{\Sigma}^{k}=K_{\epsilon}\lambda_{\Sigma}^{k}.$
for some finite $K_{\epsilon}$. Then as a direct consequence of Lemma 12, $\|Q_{k}(\theta)\|_{2}\to 0$, geometrically as $k\to\infty$ and uniformly in $\theta\in\Theta$.
We have that
$\Sigma_{\epsilon}(\theta)^{-1}\Delta\Sigma_{\epsilon,k}(\theta)=-\Sigma_{\epsilon}(\theta)^{-1}\Sigma_{\epsilon,k}(\theta)Q_{k}(\theta)\Sigma_{\epsilon}(\theta)$
Notice that $\Sigma_{\epsilon}(\theta)$ and $\Sigma_{\epsilon,k}(\theta)$ are uniformly bounded in $k$ and $\theta\in\Theta$. Since, from Lemma 13, $\|\Sigma_{\epsilon}(\theta)^{-1}\|_{2}$ is uniformly bounded for $\theta\in\Theta$, we also have $\|\Sigma_{\epsilon}(\theta)^{-1}\Delta\Sigma_{\epsilon,k}(\theta)\|_{2}\rightarrow 0$, geometrically as $k\to\infty$ and uniformly in $\theta\in\Theta$.
From $\Delta K_{k}(\theta)=K(\theta)-K_{k}(\theta)$, we have
$\Delta K_{k}(\theta)$ $=\bigl{(}F_{c}\Sigma H_{o}^{T}+S\bigr{)}\bigl{(}H_{o}\Sigma H_{o}^{T}+R\bigr{)}^{-1}$
$\quad-\bigl{(}F_{c}\Sigma_{k}H_{o}^{T}+S\bigr{)}\bigl{(}H_{o}\Sigma_{k}H_{o}^{T}+R\bigr{)}^{-1}.$
A straightforward rearrangement yields
$\Delta K_{k}=(F_{c}(\Sigma-\Sigma_{k})H_{o}^{T})\bigl{(}H_{o}\,\Sigma\,H_{o}^{T}+R\bigr{)}^{-1}$
$+\!\bigl{(}F_{c}\Sigma_{k}H_{o}^{T}\!+S\bigr{)}\!\Bigl{[}\bigl{(}H_{o}\Sigma H_{o}^{T}+R\bigr{)}^{-1}\!\!-\bigl{(}H_{o}\Sigma_{k}H_{o}^{T}+R\bigr{)}^{-1}\Bigr{]}.$
Using again the identity $A^{-1}-B^{-1}=A^{-1}(B-A)B^{-1}$, we get
$\bigl{(}H_{o}\Sigma H_{o}^{T}+R\bigr{)}^{-1}-\bigl{(}H_{o}\Sigma_{k}H_{o}^{T}+R\bigr{)}^{-1}=$
$\quad\quad\bigl{(}H_{o}\Sigma H_{o}^{T}+R\bigr{)}^{-1}H_{o}\bigl{(}\Sigma_{k}-\Sigma\bigr{)}H_{o}^{T}\bigl{(}H_{o}\Sigma_{k}H_{o}^{T}+R\bigr{)}^{-1}.$
Using the triangular inequality, we have
$\|\Delta K_{k}(\theta)\|_{2}\leq\|F_{c}\|_{2}\|\Sigma-\Sigma_{k}\|_{2}\|H_{o}\|_{2}\|(H_{o}\Sigma H_{o}^{T}+R)^{-1}\|_{2}$
$+\|F_{c}\Sigma_{k}H_{o}^{T}+S\|_{2}\|(H_{o}\Sigma H_{o}^{T}+R)^{-1}\|_{2}\|H_{o}\|_{2}\|\Sigma-\Sigma_{k}\|_{2}$
$\times\|H_{o}\|_{2}\|\bigl{(}H_{o}\Sigma_{k}H_{o}^{T}+R\bigr{)}^{-1}\|_{2}.$
Taking into account the boundedness of $\Sigma_{k}(\theta)$, $\Sigma(\theta)$, and their respective inverses, there are finite constants $M>0$ and $M_{\Sigma}>0$ such that
$\|\Delta K_{k}(\theta)\|_{2}\leq M\|\Sigma(\theta)-\Sigma_{k}(\theta)\|_{2}\leq M_{\Sigma}\lambda_{\Sigma}^{k}.$
Then, since $\|\Sigma_{k}(\theta)-\Sigma(\theta)\|$ decays with rate $\lambda_{\Sigma}$, where $0<\lambda_{\Sigma}<1$, we conclude $\|K(\theta)-K_{k}(\theta)\|$ also decays with the same rate $\lambda_{\Sigma}$. ∎
Lemma 14
Assume that A2–A3 hold uniformly in $\theta\in\Theta$. Initialize the Riccati recursion in (21) with $\Sigma_{1}(\theta)=P(\theta)$, where $P(\theta)$ is the solution of the algebraic Lyapunov equation in (22). Then it holds for the sequence of $\Sigma_{k}(\theta)$ defined by the Riccati recursion that
$\max_{i}|\lambda_{i}(F_{c}(\theta)-K_{k}(\theta)H_{o})|<\lambda_{\max}<1$
uniformly in $k\geq 1$ and $\theta\in\Theta$.
Proof:
Similarly as in (9) the Riccati recursion can be written
$\Sigma_{k+1}=\bar{F}_{c}\Sigma_{k}\bar{F}_{c}^{T}+\bar{Q}-\bar{F}_{c}\Sigma_{k}H_{o}^{T}(H_{o}\Sigma_{k}H_{o}^{T}+R)^{-1}H_{o}\Sigma_{k}\bar{F}_{c}^{T}$
Define $Q_{1}=\bar{Q}+\Sigma_{1}-\Sigma_{2}$. From (22) $\Sigma_{1}$ satisfies
$\Sigma_{1}=F_{c}\Sigma_{1}F_{c}^{T}+Q$
and we obtain from the Riccati recursion above that
$Q_{1}=Q+F_{c}\Sigma_{1}H_{o}^{T}\left(H_{o}\Sigma_{1}H_{o}^{T}+R\right)^{-1}H_{o}\Sigma_{1}F_{c}^{T}$
This implies that $Q_{1}\succeq Q$ and since $\bar{Q}=Q-SR^{-1}S^{T}$ we also have that $Q_{1}\succeq\bar{Q}$. It follows, from Lemma 10, that $(H_{o},\bar{F}_{c})$ is detectable and $(\bar{F}_{c},\bar{Q}^{1/2})$ is stabilizable, and then from Theorem 1 in [2] we have $\max_{i}|\lambda_{i}(F_{c}-K_{k}H_{o})|<1$, for all $k\geq 1$. Since $K_{k}\rightarrow K$, which satisfies the algebraic Riccati equation, it follows that the result also holds in the limit by Lemma 1. ∎
XII-B Convergence of Quadratic Form
Lemma 15
With
$\mathcal{L}_{N}(\theta)=\frac{1}{N}\sum_{k=1}^{N}\epsilon_{k}(\theta)^{T}\Sigma_{\epsilon}(\theta)^{-1}\epsilon_{k}(\theta)$
it holds w.p. 1 that
$\sup_{\theta\in\Theta}\left|\mathcal{L}_{N}(\theta)-\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta)^{T}\Sigma_{\epsilon}(\theta)^{-1}\epsilon_{1}(\theta)\,\right]\right|\rightarrow 0,\quad N\rightarrow\infty.$ (40)
Proof:
Since $\epsilon_{k}(\theta)$ is a stationary random process, $\Sigma_{\epsilon}(\theta)\succ 0$ by Lemma 1, and since
$\operatorname{\mathbb{E}}\!\left[\,\left|\epsilon_{1}(\theta)^{T}\Sigma_{\epsilon}(\theta)^{-1}\epsilon_{1}(\theta)\right|\,\right]=1<\infty,$
it holds by the strong law of large numbers, see [6, Theorem 5.4.2], that we have point-wise convergence for every $\theta\in\Theta$. To prove uniform convergence we may, since $\Theta$ is a compact set, show this by first showing that $\left\|\frac{\partial\mathcal{L}_{N}}{\partial\theta_{i}}\right\|_{2}$ is bounded w.p. 1 and uniformly in $\theta\in\Theta$ and in $N\geq 1$. Then we can use the mean value theorem, from which the lemma follows. For these latter details we refer the reader to [4]. Here we will prove the boundedness of the partial derivatives. It holds that
$\frac{\partial\mathcal{L}_{N}}{\partial\theta_{i}}$ $=\frac{1}{N}\sum_{k=1}^{N}\frac{\partial\epsilon_{k}(\theta)^{T}}{\partial\theta_{i}}\Sigma_{\epsilon}(\theta)^{-1}\epsilon_{k}(\theta)$ (41)
$-\frac{1}{N}\sum_{k=1}^{N}\epsilon_{k}(\theta)^{T}\Sigma_{\epsilon}(\theta)^{-1}\frac{\partial\Sigma_{\epsilon}(\theta)}{\partial\theta_{i}}\Sigma_{\epsilon}(\theta)^{-1}\epsilon_{k}(\theta)$
$+\frac{1}{N}\sum_{k=1}^{N}\epsilon_{k}(\theta)^{T}\Sigma_{\epsilon}(\theta)^{-1}\frac{\partial\epsilon_{k}(\theta)}{\partial\theta_{i}}.$
From now on we will denote derivative with respect to $\theta_{i}$ with a prime, i.e. $\epsilon^{\prime}_{k}(\theta)=\frac{\partial\epsilon_{k}(\theta)}{\partial\theta_{i}}$, and we will also sometimes omit the dependence on $\theta$. From (4) it follows that
$\hat{\xi}^{\prime}_{k+1}$ $=\left(F_{c}-KH_{o}\right)\hat{\xi}^{\prime}_{k}+F_{c}^{\prime}\hat{\xi}_{k}+K^{\prime}\epsilon_{k}$
$\epsilon^{\prime}_{k}$ $=-H_{o}\hat{\xi}^{\prime}_{k},$
where $\hat{\xi}_{k}$ is defined in (5) and and $\epsilon_{k}=x_{o,k}-H_{o}\hat{\xi}_{k}-J_{ro}r_{k}$. The latter is a stationary stochastic process uniformly in $\theta\in\Theta$. From Lemma 5 it follows that the solution $(\Sigma,K,\Sigma_{\epsilon})$ of the algebraic Riccati equation in (3) is continuously differentiable for every $\theta\in\Theta_{o}$. This also holds for $F_{c}$. Notice that $\hat{\xi}_{k}$ is quasi-stationary. Hence it follows that $\epsilon^{\prime}_{k}$ is a quasi-stationary stochastic processes uniformly in $\theta\in\Theta$ if $r_{k}$ is quasi-stationary. Therefore there exist $\rho<1$ and $\kappa_{i}$ independent of $\theta$ such that
$\|\epsilon_{k}(\theta)\|_{2}$ $\leq\kappa_{1}\sum_{j=-\infty}^{k}\rho^{k-j}\|x_{o,k}\|_{2}+\kappa_{2}\sum_{j=-\infty}^{k}\rho^{k-j}\|r_{k}\|_{2}$
$\|\epsilon_{k}^{\prime}(\theta)\|_{2}$ $\leq\kappa_{3}\sum_{j=-\infty}^{k}\rho^{k-j}\|x_{o,k}\|_{2}+\kappa_{4}\sum_{j=-\infty}^{k}\rho^{k-j}\|r_{k}\|_{2}.$
From this it follows for some $\kappa_{5}$ independent of $\theta$, since $\Sigma_{\epsilon}\succ 0$ and its derivative is continuously differentiable for every $\theta\in\Theta$ that
$\left\|\frac{\partial\mathcal{L}_{N}}{\partial\theta_{i}}\right\|_{2}\!\!\leq\!\frac{\kappa_{5}}{N}\sum_{k=1}^{N}\!\left(\sum_{j=-\infty}^{k}\!\!\rho^{k-j}\|x_{o,k}\|_{2}+\!\!\!\sum_{j=-\infty}^{k}\rho^{k-j}\|r_{k}\|_{2}\!\right)^{2}.$
By the strong law of large numbers the right-hand side converges w.p. 1 to
$\kappa_{5}\operatorname{\mathbb{E}}\!\left[\,\left(\sum_{j=-\infty}^{k}\rho^{k-j}\|x_{o,k}\|_{2}+\sum_{j=-\infty}^{k}\rho^{k-j}\|r_{k}\|_{2}\right)^{2}\,\right]$
if this expectation is finite, which is straightforward to show. Hence we have that $\left\|\frac{\partial\mathcal{L}_{N}}{\partial\theta_{i}}\right\|_{2}$ is bounded w.p. 1 and uniformly in $\theta\in\Theta$ and in $N\geq 1$. ∎
XII-C Convergence of Log-Likelihood Function
We have the following lemma:
Lemma 16
Assume that A2-A3 hold uniformly in $\theta\in\Theta$. It then holds that
$\sup_{\theta\in\Theta}\left|L_{N}(\theta)-L(\theta)\right|\rightarrow 0,\quad N\rightarrow\infty\;\mathrm{w.p.1}$
where
$L(\theta)=-\operatorname{\mathbb{E}}\!\left[\,\ln f(\epsilon_{1}(\theta);\theta)\,\right]$ (42)
and where
$f(\epsilon;\theta)=\frac{1}{\sqrt{(2\pi)^{n_{0}}\det\Sigma_{\epsilon}(\theta)}}\exp\left(-\frac{1}{2}\epsilon^{T}\Sigma_{\epsilon}(\theta)^{-1}\epsilon\right).$
Proof:
We follow the lines of the proof in [20]. Let
$\Delta^{\prime}_{N}(\theta)=\frac{1}{N}\sum_{k=1}^{N}\left(\ln\det\Sigma_{\epsilon,k}(\theta)-\ln\det\Sigma_{\epsilon}(\theta)\right),\quad N\geq 1$
and
$\Delta^{\prime\prime}_{N}(\theta)=\frac{1}{N}\sum_{k=1}^{N}\bigg{(}$ $\check{\epsilon}_{k}(\theta)^{T}\Sigma_{\epsilon,k}(\theta)^{-1}\check{\epsilon}_{k}(\theta)$
$-\epsilon_{k}(\theta)^{T}\Sigma_{\epsilon}(\theta)^{-1}\epsilon_{k}(\theta)\bigg{)}.$
From $\Sigma_{\epsilon,k}(\theta)=\Sigma_{\epsilon}(\theta)+\Delta\Sigma_{\epsilon,k}(\theta)=\Sigma_{\epsilon}(\theta)\left(I+\Sigma_{\epsilon}(\theta)^{-1}\Delta\Sigma_{\epsilon,k}(\theta)\right)$, with obvious definition of $\Delta\Sigma_{\epsilon,k}(\theta)$, we have that
$\Delta^{\prime}_{N}(\theta)=\frac{1}{N}\sum_{k=1}^{N}\ln\det\left(I+\Sigma_{\epsilon}(\theta)^{-1}\Delta\Sigma_{\epsilon,k}(\theta)\right).$
By Corollary 13.1 it holds that $\|\Sigma_{\epsilon}(\theta)^{-1}\Delta\Sigma_{\epsilon,k}(\theta)\|\rightarrow 0$ both uniformly and geometrically as $\rightarrow\infty$. Hence the eigenvalues of $I+\Sigma_{\epsilon}(\theta)^{-1}\Delta\Sigma_{\epsilon,k}(\theta)$ converges to $1$ uniformly in $\theta\in\Theta$ and geometrically in $k$ as $k\rightarrow\infty$. Hence we have that
$\sup_{\theta\in\Theta}\left|\Delta^{\prime}_{N}(\theta)\right|\rightarrow 0,\quad N\rightarrow\infty.$
We now focus on $\Delta^{\prime\prime}_{N}(\theta)$. Let
$d_{k}(\theta)$ $=\check{\epsilon}_{k}(\theta)-\epsilon_{k}(\theta)$
$Q_{k}(\theta)$ $=\Sigma_{\epsilon,k}(\theta)^{-1}-\Sigma_{\epsilon}(\theta)^{-1}.$
We can then write
$\Delta^{\prime\prime}_{N}(\theta)=\frac{1}{N}\sum_{k=1}^{N}\bigg{(}\epsilon_{k}(\theta)^{T}Q_{k}(\theta)\epsilon_{k}(\theta)$ (43)
$\quad\quad+2d_{k}(\theta)^{T}\Sigma_{\epsilon,k}(\theta)^{-1}\epsilon_{k}(\theta)+d_{k}(\theta)^{T}\Sigma_{\epsilon,k}(\theta)^{-1}d_{k}(\theta)\bigg{)}.$
By Corollary 13.1 $\|Q_{k}(\theta)\|\rightarrow 0$ geometrically as $k\rightarrow\infty$ and uniformly in $\theta\in\Theta$, and by the strong law of large numbers, [6, Theorem 5.4.2].
[…]
The eigenvalues of $F_{c}(\theta)$ are uniformly strictly inside the unit circle, and hence $\operatorname{\mathbb{E}}\!\left[\,\epsilon_{1}(\theta)^{T}\epsilon_{1}(\theta)\,\right]$ is uniformly bounded with respect to $\theta\in\Theta$. Hence the first sum in (43) converges to zero w.p. 1 and uniformly in $\theta\in\Theta$.
By Lemma 13 we have that $\Sigma_{\epsilon,k}(\theta)^{-1}<\kappa_{1}I$ for some positive $\kappa_{1}$. Hence the absolute value of the second sum in (43) is by the Cauchy-Schwartz inequality majorized by
$2\kappa_{1}\left(\frac{1}{N}\sum_{k=1}^{N}d_{k}(\theta)^{T}d_{k}(\theta)\right)^{1/2}\left(\frac{1}{N}\sum_{k=1}^{N}\epsilon_{k}(\theta)^{T}\epsilon_{k}(\theta)\right)^{1/2}$
and the third sum by
$\frac{\kappa_{1}}{N}\sum_{k=1}^{N}d_{k}(\theta)^{T}d_{k}(\theta).$
Therefore, in order to prove that the second and third sum converges to zero, it is sufficient to prove that
$\frac{1}{N}\sum_{k=1}^{N}d_{k}(\theta)^{T}d_{k}(\theta)\rightarrow 0,\quad N\rightarrow\infty\;\mathrm{w.p.1}$ (44)
and uniformly in $\theta\in\Theta$. From the definitions of $d_{k}(\theta)$, $\epsilon_{k}(\theta)$ and $\check{\epsilon}_{k}(\theta)$ it follows with $\tilde{\xi}_{k}(\theta)=\hat{\xi}_{k}(\theta)-\check{\xi}_{k}(\theta)$ that
$\tilde{\xi}_{k+1}(\theta)$ $=\tilde{F}_{k}(\theta)\tilde{\xi}_{k}+\Delta K_{k}(\theta)\epsilon_{k}(\theta)$
$d_{k}(\theta)$ $=H_{o}\tilde{\xi}_{k}(\theta),$
where $\tilde{F}_{k}(\theta)=F_{c}(\theta)-K_{k}(\theta)H_{o}$ and $\Delta K_{k}(\theta)=K(\theta)-K_{k}(\theta)$. The initial value is $\tilde{\xi}_{1}(\theta)=\hat{\xi}_{1}(\theta)$. Define the transition matrix
$\Phi_{k,j}(\theta)=\tilde{F}_{k-1}(\theta)\tilde{F}_{k-2}(\theta)\cdots\tilde{F}_{j}(\theta).$
Then we can write the solution to the difference equation above as $d_{1}(\theta)=H_{o}\hat{\xi}_{1}(\theta)$ and for $k\geq 2$
$d_{k}(\theta)=H_{o}\Phi_{k,1}(\theta)\hat{\xi}_{1}(\theta)+\sum_{j=1}^{k-1}H_{o}\Phi_{k,j+1}(\theta)\Delta K_{j}(\theta)\epsilon_{j}(\theta).$
Let $\alpha_{1}=H_{o}\hat{\xi}_{1}(\theta)$ and $\beta_{1}=0$ and for $k\geq 2$
$\alpha_{k}$ $=H_{o}\Phi_{k,1}(\theta)\hat{\xi}_{1}(\theta)$
$\beta_{k}$ $=\sum_{j=1}^{k-1}H_{o}\Phi_{k,j+1}(\theta)\Delta K_{j}(\theta)\epsilon_{j}(\theta).$
We obtain from the Cauchy-Schwartz inequality that
$\frac{1}{N}\sum_{k=1}^{N}d_{k}(\theta)^{T}d_{k}(\theta)\leq$
$\frac{1}{N}\!\Bigg{[}\sum_{k=1}^{N}\alpha_{k}^{T}\alpha_{k}+2\Bigg{(}\!\sum_{k=1}^{N}\alpha_{k}^{T}\alpha_{k}\Bigg{)}^{\frac{1}{2}}\!\!\Bigg{(}\!\sum_{k=1}^{N}\beta_{k}^{T}\beta_{k}\Bigg{)}^{\frac{1}{2}}\!\!\!\!+\!\sum_{k=1}^{N}\beta_{k}^{T}\beta_{k}\Bigg{]}.$
It is now sufficient to show that
$\frac{1}{N}\sum_{k=1}^{N}\alpha_{k}^{T}\alpha_{k}\rightarrow 0,\quad\frac{1}{N}\sum_{k=1}^{N}\beta_{k}^{T}\beta_{k}\rightarrow 0$
as $N\rightarrow\infty$ w.p. 1 and uniformly in $\theta\in\Theta$.
It holds by Lemma 14 that $\max|\lambda_{i}(\tilde{F}_{k}(\theta))|<\lambda_{\max}<1$ uniformly in $k\geq 1$ and $\theta\in\Theta$. Hence
$\|\Phi_{k,j}(\theta)\|\leq\lambda_{\max}^{k-j}$
and we obtain
$\frac{1}{N}\sum_{k=1}^{N}\alpha_{k}^{T}\alpha_{k}\leq\frac{\|H_{o}\|^{2}\|\hat{\xi}_{1}(\theta)\|_{2}^{2}}{N}\sum_{k=1}^{N}\lambda_{\max}^{2(k-1)}\rightarrow 0,\quad N\rightarrow\infty$
w.p. 1 uniformly in $\theta\in\Theta$.
By Corollary 13.1 it hold that $\|\Delta K_{k}(\theta)\|<M\kappa_{2}^{k}$ where $\kappa_{2}<1$ uniformly in $\theta\in\Theta$ and where $M$ is a positive constant. If $\lambda_{\max}\leq\kappa_{2}$ we can always find a new value of $\lambda_{max}<1$ that is greater than $\kappa_{2}$. Hence we will without loss in generality assume that $\kappa_{3}=\kappa_{2}/\lambda_{\max}<1$. We then obtain that
$\frac{1}{N}\sum_{k=1}^{N}\beta_{k}^{T}\beta_{k}$ $\leq\frac{M^{2}\|H_{o}\|^{2}}{N}\sum_{k=2}^{N}\sum_{i=1}^{k-1}\lambda_{\max}^{k-i-1}\kappa_{2}^{i}\|\epsilon_{i}\|_{2}$
$\times\sum_{j=1}^{k-1}\lambda_{\max}^{k-j-1}\kappa_{2}^{j}\|\epsilon_{j}\|_{2}$
$<\frac{M^{2}\|H_{o}\|^{2}}{N}\sum_{k=2}^{N}\lambda_{\max}^{2(k-1)}\sum_{i,j=1}^{k-1}\kappa_{3}^{i+j}\|\epsilon_{i}\|_{2}\|\epsilon_{j}\|_{2}$
$<\frac{M^{2}\|H_{o}\|^{2}}{N}\sum_{k=2}^{N}(k-1)^{2}\lambda_{\max}^{2(k-1)}\frac{1}{(k-1)^{2}}$
$\times\sum_{i,j=1}^{k-1}\|\epsilon_{i}\|_{2}\|\epsilon_{j}\|_{2}.$
As before, since $\epsilon_{j}$ is a stationary stochastic process, it holds by the strong law of large numbers that $\frac{1}{k-1}\sum_{j=1}^{k-1}\|\epsilon_{j}\|_{2}$ has a finite limit w.p. 1 as $k\rightarrow\infty$. This convergence is uniform in $\theta\in\Theta$ with a similar argument as in the previous section. Hence the sum above is uniformly bounded in $k\geq 1$ and $\theta\in\Theta$ w.p. 1. From this we obtain that
$\frac{1}{N}\sum_{k=1}^{N}\beta_{k}^{T}\beta_{k}\rightarrow 0,\quad N\rightarrow\infty\;\mathrm{w.p.1}$
uniformly in $\theta\in\Theta$.
We have now proven that both $\Delta^{\prime}N(\theta)$ and $\Delta^{\prime\prime}N(\theta)$ converges to zero as $N$ goes to infinity uniformly in $\theta\in\theta$, and hence by Lemma 15 we have proven this lemma. ∎